### v2

# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

Abstract
Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train a deep learning model by submitting hidden model updates. Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients. Existing solutions have struggled to improve the performance of cross-silo PPFL in poisoned Non-IID data. To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while effectively resisting data poisoning attacks. Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of tampered data distribution on federated learning. Moreover, we utilize two servers to achieve Byzantine-robust aggregation by secure aggregation protocol, which greatly reduces the impact of malicious clients. Theoretical analyses confirm the convergence of PPFPL, and experimental results on publicly available datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID conditions.
keywords:
Privacy-preserving, cross-silo, federated learning, data poisoning attacks, poisoned Non-IID data.
journal: Nuclear Physics B
1 Introduction
Federated Learning (FL) is a distributed learning paradigm where each client submits its model updates instead of raw training data. In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1]. To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3]. Specifically, most clients in cross-silo FL are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities. However, client-submitted model updates are vulnerable to privacy inference attacks, which threats the privacy security of cross-silo FL [4][5]. To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] or Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed. In particular, the DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the accuracy of the global model by introducing noise. Conversely, the HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
While HE-based works [8] [9] demonstrated their effectiveness in privacy preservation within cross-silo FL, they are susceptible to data poisoning attacks due to the distributed training paradigm [10][11]. Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, severely impacting the performance of PPFL [12][13]. Moreover, in PPFL, data poisoning attacks are more concealed due to privacy technology obscuring model updates from malicious clients, which raises the difficulty of defense. To audit obscured model updates, recent works [14][15][16][17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious updates in ciphertext. However, these works overlook the heterogeneity of local training data (i.e., Non-IID) issue in FL, which causes distinct inconsistency of benign updates, thereby disrupting the defense against malicious updates.
To distinguish between benign and malicious updates in the Non-IID data, existing solutions [18][19] incorporate clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase. While these solutions are effective against model poisoning attacks, they are inadequate for countering data poisoning attacks. This limitation stems from their reliance on specialized training techniques designed for Non-IID data (e.g., FedProx[20], FedDyn[21], or FedLC[22], etc.), which enhance the performance of global model in Non-IID scenarios. However, in data poisoning attacks, malicious clients deliberately tamper with the features and labels of their training data, thereby forming poisoned Non-IID data. These manipulations distort the optimization direction of local model training and significantly degrade the effectiveness of the specialized training techniques. Consequently, integrating the above defense solutions [18][19] with such training techniques fails to effectively mitigate data poisoning attacks in poisoned Non-IID data. Therefore, a critical challenge remains: how to improve the performance of cross-silo PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
Inspired by prototype learning, several works [23][24] designed prototype aggregation to address the Non-IID issue in federated learning, where prototypes are transmitted between the server and clients. Each prototype represents a class feature that is calculated from the mean representations transformed from the samples belonging to the same class. For instance, the works in [23][24] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”. By exchanging these prototypes, clients can gain more knowledge about the concept of “cat”. Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
In cases of poisoned Non-IID data, tampered data features and data distribution cannot be repaired since the server cannot manipulate the behavior of malicious clients. This prompts us to pose a question: Is it possible to design a PPFL that leverages prototype learning, so that client-submitted model updates are affected only by tampered data features and not by tampered data distribution, while incorporating a secure aggregation protocol to achieve Byzantine-robust results? To answer this question, we propose Privacy-Preserving Federated Prototype Learning framework, named PPFPL, which is suitable for cross-silo scenarios. Our core idea is to utilize prototype learning to address the issue of tampered data distribution, while designing a secure aggregation protocol across two servers to achieve Byzantine-robust aggregation results without compromising privacy. PPFPL employs two “non-colluding” servers who are “honest but curious” to execute a secure aggregation protocol that utilizes cosine similarity as a method to penalize malicious clients. In addition, HE and SMC are adopted as underlying technologies to add the privacy-preserving mechanism of the secure aggregation protocol. The main contributions of the paper are summarized as follows.
  • 1.
    We are the first to introduce prototype learning into PPFL to address data poisoning attacks in across-silo FL, which eliminates the impact of tampered data distribution for client-submitted model updates while significantly improving the FL performance in poisoned Non-IID data.
  • 2.
    We design a secure aggregation protocol across two servers to aggregate client-submitted prototypes with HE and SMC techniques, which ensures that two servers and malicious clients cannot utilize client-submitted prototypes to invade the privacy of benign clients.
  • 3.
    We provide a convergence guarantee for PPFPL in the presence of data poisoning attacks, which theoretically ensures the framework’s feasibility.
  • 4.
    Compared to existing methods, the superiority of our framework has been empirically validated in poisoned Non-IID data.
The rest of the paper is organized as follows. Section 2 reviews recent PPFL works against poisoning attacks. Section 3 introduces prototype learning and HE. Section 4 formalizes the system model of PPFPL. Our framework is presented in Section 5. Section 6 provide theoretical analysis. Section 7 reviews experimental results. Finally, Section 8 concludes this paper.
2 RELATED WORK
Privacy preservation in FL: Although across-silo FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of clients. To resist privacy attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates. Specifically, DP-based schemes [6][7][25][26] deploy local differential privacy into model updates to ensure privacy without compromising the utility of model updates. Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which reduces the performance of their global models to some extent. In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without degrading the global model’s accuracy. Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [8]. They proposed a multi-party machine learning scheme using Paillier [27] technique without compromising participants’ private data. Considering the heavy communication overhead of Paillier, the work in [9] proposed a privacy-preserving federated learning scheme using CKKS (i.e., Cheon-Kim-Kim-Song). This scheme significantly reduces computational overhead associated with ciphertexts. This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [28]. However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
Resisting data poisoning attacks in FL: Tolpegin et al. demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [29]. Additionally, they proposed an aggregated defense strategy that can identify malicious clients in FL to circumvent data poisoning attacks. In addition, Zhang et al. proposed an effective framework capable of tracking the attack time, objective, type, and poisoned location for model updates [30]. Differently, Doku et al. employed SVM to audit client’s local training data for excluding malicious clients, which prevents tampered data from poisoning the global model [31]. However, this violates the privacy of clients to some extent. Furthermore, in practice, the threats of privacy and data poisoning attacks usually coexist in across-silo FL. PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while also concealing data poisoning attacks from malicious clients.
Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15] [14][16][17] combine HE and SMC to identify anomalous model updates in ciphertexts. For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function. Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks. However, these schemes [15] [14][16][17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious updates or Non-IID data.
To audit malicious updates in PPFL with Non-IID data, the works in [18][19] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage. Specifically, ShieldFL designs a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [18]. Furthermore, Chen et al. adopts clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [19]. These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data. They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[20], FedDyn[21], or FedLC[22], etc.) to radically improve performance of FL. These specialized techniques adds an auxiliary term into local loss function to constrain model updates of clients, which helps to increase consistency of model updates during local model training. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs. Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while improving the performance of the global model under poisoned Non-IID data.
Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue. Specifically, the works in [23][24][32] are one of the first to propose federated prototype learning using the concept of prototype learning. Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), its core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes. In other words, each class holds its corresponding prototype that is independent of other classes. As a result, client-submitted prototype is affected by the samples and is independent of data distribution among clients. This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks. Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation in across-silo FL caused by tampered Non-IID data, while preserving the privacy of client-submitted prototypes.
3 PRELIMINARIES
This section introduces prototypes in federated learning and CKKS technology.
3.1 Prototypes Meet Federated Learning
In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [33]. Due to this property, the prototypes of the same class among clients are similar in FL task. Consequently, many FL schemes [23][32] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
To further understand the prototype calculation in federated prototype learning, we introduce some basic notations below. Let $\mathcal{S}$ be the set of clients, where each client $m\in\mathcal{S}$ has an independent private dataset, denoted as $\mathcal{D}_{m}=\{(\textit{{x}}_{(i)},y_{(i)})\}^{|\mathcal{D}_{m}|}$. Here, $|\mathcal{D}_{m}|$ represents the number of samples in client $m$, and $(\textit{{x}}_{(i)},y_{(i)})$ denotes sample $i$ in dataset, where $\textit{{x}}_{(i)}$ and $y_{(i)}$ correspond to the feature vector and class label of sample $i$, respectively. Meanwhile, let $\mathcal{I}$ be the set of classes in classification task, where each class $k$ belongs to $\mathcal{I}$. In classification task, the local model includes a feature extractor and a decision classifier. Specifically, the feature extractor transforms sample features into compressed features, while decision classifier maps the compressed features to get classification results. Formally, let $f_{m}(\textit{{r}}_{m,t};\cdot)$ be feature extractor for client $m$, parameterized by $\textit{{r}}_{m,t}$, where $t$ denotes the $t$-th communication round. Given the feature $\textit{{x}}_{(i)}$ of sample $i$, it is input to feature extractor to obtain compressed feature $\textit{{u}}_{(i)}=f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)})$. Let $g_{m}(\textit{{z}}_{m,t};\cdot)$ be decision classifier for client $m$, parameterized by $\textit{{z}}_{m,t}$. The classifier maps the compressed feature $\textit{{u}}_{(i)}$ to predict the class $y^{\prime}=g_{m}(\textit{{z}}_{m,t};\textit{{u}}_{(i)})$. Thus, we denote the local model as $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=g_{m}(\textit{{z}}_{m,t};\cdot)\circ f_{m}(\textit{{r}}_{m,t};\cdot)$, where $\circ$ denotes composite operator. To simplify notation, we use $\textit{{w}}_{m,t}$ to denote $(\textit{{r}}_{m,t},\textit{{z}}_{m,t})$, so we have $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=\mathcal{F}_{m}(\textit{{w}}_{m,t};\cdot)$, and $\textit{{w}}_{m,t}$ is consider as model parameters for client $m$. Next, we introduce calculation process of prototypes.
In federated prototype learning, prototypes can be categorized into local prototypes, computed by clients, and global prototypes, aggregated by the server. Specifically, each client’s goal is to align its local prototype to global prototype during local model training. Thus, each client computes its local prototype via its training dataset during local model training. Formally, let $\textit{{c}}^{k}_{m,t}$ be the local prototype of class $k\in\mathcal{I}$ at client $m$ in $t$-th communication round, calculated as
$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$ (1)
where $\mathcal{D}^{k}_{m}$ denotes the dataset with class $k$ at client $m$. The $\textit{{c}}^{k}_{m,t}$ can be understood as the mean of compressed features of samples belonging to class $k$ at client $m$. Further, we can get local prototypes for all classes at client $m$, denoted as $\{\textit{{c}}^{k}_{m,t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$. Once local model training is complete, each client submits its local prototypes to server.
To calculate global prototype, the server adopts an averaging operation to local prototypes submitted by clients. Thus, the global prototype $\textit{{C}}^{k}_{t+1}$ for each class is calculated as follows:
$\textit{{C}}^{k}_{t+1}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\textit{{c}}_{m,t}^{k},\forall k\in\mathcal{I},$ (2)
where $|\mathcal{S}|$ denotes the number of clients. For global prototypes of all classes, we denote them by the set $\{\textit{{C}}^{k}_{t+1}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$. Subsequently, the server distributes the latest global prototypes to each client to further train their local model.
3.2 CKKS
CKKS is a homomorphic encryption technology. Similar to traditional HE, it is a cryptographic technique based on mathematical computations, allowing operations on encrypted data without decrypting. However, traditional HE schemes such as RSA [34], ElGamal [35], and Paillier are limited in that they only support either additive or multiplicative operations but not both simultaneously. In contrast, the CKKS technology provides both additive and multiplicative homomorphic encryption, which is known as full HE [36]. In addition, CKKS is known for its efficiency, especially in terms of encryption/decryption speed when dealing with scale vectors of varying parameter lengths. Therefore, we employ CKKS for the privacy protection of clients due to its computational efficiency. For a more detailed implementation principle and process of CKKS, please refer to paper [37].
4 Problem Statement
In the section, we formalize the PPFPL framework, define potential threats, and design goals.
4.1 PPFPL framework
The framework of PPFPL consists of four entities, each of which has its specific function, as shown in Fig. 1. The interaction among entities composes operation of whole system. The specific functions of each entity are outlined as follows.
Fig. 1: The PPFPL framework. ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients. ② After local training is completed, clients submit encrypted local prototypes to the Aggregator. ③ Verifier and Aggregator perform secure aggregation protocol to get encrypted global prototypes. ④ Aggregator distributes encrypted global prototypes to Clients.
  • 1.
    Key Generation Center (KGC). The entity is responsible for generating and managing keys of both Clients and Verifier, which are important elements to ensure security of encryption/decryption process.
  • 2.
    Clients. The Clients are large organizations participating in federated training. The aim of benign organizations is to get a better model by federated training. They has a pair of public/secret keys generated by KGC, denoted as $Pkx$/$Skx$.
  • 3.
    Aggregator. The Aggregator is a central server responsible for aggregating local prototypes submitted by clients.
  • 4.
    Verifier. The Verifier is a non-colluding central server and calculates cooperatively with Aggregator to aggregate local prototypes. It has a pair of public/private keys generated by KGC, denoted as $Pkv$/$Skv$.
In our framework, we not only define the function of each entity, but also define potential threats.
4.2 Potential Threats for PPFPL
We discuss potential threats of PPFPL in detail.
i) The KGC is a trusted institution (e.g., government, union).
ii) The Aggregator and Verifier are considered as “non-colluding” and “curious but honest”. Specifically, we assume that Aggregator and Verifier do not collude to attack PPFPL. The assumption is reasonable in practice, since it is usually impossible for two well-known service providers to collude with each other due to legal regulations, company reputation [38]. Furthermore, we assume that they follow the system’s protocol but may attempt to get sensitive information (i.e., raw training data) by inferring client-submitted local prototypes. Notably, although prototype is different from gradient, traditional inference attack methods cannot recover client’s private training data from prototype. However, the work in [39] designs a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes. Thus, preserving the privacy of prototypes is necessary. In addition, we do not consider security attacks resulting from employee insider threats (e.g., compromised employees within the Aggregator or Verifier).
iii) In practice, we cannot guarantee that all clients in the system are honest. Therefore, clients can be either benign or malicious. Specifically, benign clients are “honest but curious”. For malicious clients, they can collude together to infer sensitive information about benign clients. We consider that the proportion of malicious clients is less than 50%, which is a more realistic threat in cross-silo FL due to the high reputation of large organizations participating. Furthermore, malicious clients can launch data poisoning attacks, and submit malicious local prototypes derived from the poisoned data. In federated prototype learning, we define the following two representative types of data poisoning attacks from the perspective of data features and labels.
  • 1.
    Feature attacks. The purpose of feature attacks is to degrade the performance of federated learning. Specifically, malicious clients tamper with the features of their training data to generate malicious prototypes for uploading, thereby affecting the training results of other clients.
  • 2.
    Label attacks. The malicious clients tamper with the labels of their training samples into other random labels to form the tampered data distribution, aiming to reduce the performance of the entire federated learning.
4.3 Design Goals of PPFPL
The goal of our study is to improve performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Specifically, we design the PPFPL framework to fulfill the following goals:
  • 1.
    Security. PPFPL should guarantee the correct model training in the presence of data poisoning attacks with Non-ID data. In other words, the model performance of each benign client is not affected by data poisoning attacks under different data distributions.
  • 2.
    Privacy. PPFPL should ensure privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
  • 3.
    Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar frameworks, thereby reducing privacy computation overheads and communication overheads.
5 Design of PPFPL
In this section, we first provide the overview of PPFPL, and then describe each step in detail.
5.1 Overview of PPFPL
The execution process of PPFPL is summarized in Algorithm 1. The framework initializes the number of communication rounds $T$, the number of local iterations $E$, and the local model’s parameters $\textit{{w}}^{\textit{init}}$ for each client. Then, Aggregator, Verifier, and Clients jointly perform FL training. Specifically, PPFPL iteratively performs the following two steps:
  • 1.
    Step I. Local Computation: Each client $m$ trains its local model with its local dataset $\mathcal{D}_{m}$. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
  • 2.
    Step II. Secure Aggregation Protocol: The two servers verify the normalization of encrypted local prototypes submitted by clients, and performs the secure two-party computation to get encrypted global prototypes, and distribute them to each client.
The above process is stopped until configured number of communication rounds $T$. In the following, we describe the process of Steps I and II in detail.
Input: $\mathcal{S}$, $\textit{{w}}^{\textit{init}}$, $E$, $T$.
Output: Model parameter of each client.
1 Initialize $T,E,\textit{{w}}^{\textit{init}}$;
2 Aggregator distributes $T,E,\textit{{w}}^{\textit{init}}$ to each client;
3 for each communication round $t\in\{1,2,\cdots,T\}$ do
4       // Step I: $\mathtt{Local\ Computation}$.
5       for each client $m\in\mathcal{S}$ do
6             Train local model;
7             Normalize and encrypt local prototypes;
8             Send encrypted local prototypes to Aggregator;
9
10      // Step II: $\mathtt{Secure\ Aggregation\ Protocol}$.
11       Two servers verify normalization;
12       Two servers compute global prototypes;
13       Aggregator distributes global prototypes to each client;
14
return Model parameter of each client
Algorithm 1 Overview of PPFPL
5.2 Local Computation
The local computation step includes two essential stages: local model training and prototype handling. The detail is outlined in Algorithm 2.
Input: $\mathcal{S}$, $\mathcal{D}_{m}$, $\eta$ , $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$, $E$.
Output: Encrypted local prototypes.
1 for each client $m$in $\mathcal{S}$ do
2       Get global prototypes $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ from Aggregator;
3       Decrypt $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ using its $Skx$;
4       $\textit{{w}}_{m,t}^{(E)}\leftarrow\textbf{Training}$($\textit{{w}}_{m,t-1}^{(E)}$, $\mathcal{D}_{m}$, $\eta$, $E$, $\{\textit{{C}}^{k}_{t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$);
5       $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}\leftarrow\textbf{Handling}(\textit{{w}}_{m,t}^{(E)},\mathcal{D}_{m})$;
6       Send $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator;
7
return $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
Algorithm 2 Local Computation
5.2.1 Local Model Training
During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype. To achieve this, an auxiliary term is incorporated into the local loss function. Formally, the local loss function of client $m$ is defined as:
$\mathcal{L}(\textit{{w}}_{m};\mathcal{D}$ ,${}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})$ (3)
$+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k}),\forall(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}_{m},\forall k\in\mathcal{I},$
where $\mathcal{L_{S}}(\cdot,\cdot)$ is the classification loss function (e.g., cross-entropy loss function), $\lambda$ is the importance weight of auxiliary term, and $\mathcal{L_{R}}(\cdot,\cdot)$ is the auxiliary term, defined as
$\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\frac{1}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}(1-\textit{sim}(\textit{{c}}_{m}^{k},\textit{{C}}^{k})),$ (4)
where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors. The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions. The objective of each client aims to minimize classification loss while keeping its local prototype close to the global prototype in the direction. Next, we present the process of local model training.
Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model. This iterative process is executed in Algorithm 3. The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$. Firstly, each client $m$ uses the local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round. Subsequently, each client iteratively performs the following stages.
  • 1.
    In $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
  • 2.
    The client inputs the training data $\mathcal{D}_{m}^{(e)}$ into the local model’s feature extractor to compute the local prototype $\textit{{c}}^{k,(e)}_{m,t}$ via formula (1).
  • 3.
    The client computes the unbiased stochastic gradient by
    $\textit{{g}}_{m,t}^{(e-1)}=\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e-1)};\mathcal{D}_{m}^{(e)},\textit{{c}}^{k,(e)}_{m,t},\textit{{C}}^{k}_{t}),$
    where $\nabla$ denotes derivation operation. Then, the local model $\textit{{w}}_{m,t}^{(e-1)}$ is updated via the following formula:
    $\textit{{w}}_{m,t}^{(e)}=\textit{{w}}_{m,t}^{(e-1)}-\eta\textit{{g}}_{m,t}^{(e-1)},$
    where $\eta$ is the local learning rate.
After $E$ local iterations of the above process, each client derives the updated local model parameters $\textit{{w}}_{m,t}^{(E)}$.
Input: $\textit{{w}}_{m,t-1}^{(E)}$, $\mathcal{D}_{m}$, $\eta$, $E$, $\{\textit{{C}}^{k}_{t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$.
Output: Local model parameters of each client $\textit{{w}}_{m,t}^{(E)}$.
1 $\textit{{w}}_{m,t}^{(0)}=\textit{{w}}_{m,t-1}^{(E)}$;
2 for each local iteration $e$in $\{1,2,\cdots,E\}$ do
3       Randomly sample $\mathcal{D}_{m}^{(e)}\subset\mathcal{D}_{m}$;
4       for each class $k$in $\mathcal{I}$ do
5             Calculate $\textit{{c}}^{k,(e)}_{m,t}$ from $\mathcal{D}^{k,(e)}_{m}$ with formula (1);
6      $\textit{{g}}_{m,t}^{(e-1)}=\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e-1)};\mathcal{D}_{m}^{(e)},\textit{{c}}^{k,(e)}_{m,t},\textit{{C}}^{k}_{t})$;
7       $\textit{{w}}_{m,t}^{(e)}=\textit{{w}}_{m,t}^{(e-1)}-\eta\textit{{g}}_{m,t}^{(e-1)}$;
8
return $\textit{{w}}_{m,t}^{(E)}$
Algorithm 3 Training
5.2.2 Prototype Handling
The prototype handling stage consists of three key phases: prototype generation, normalization, and encryption, as illustrated in Algorithm 4.
Prototype Generation
Since the parameters of local model change with each iteration, the prototypes generated by each local iteration are different. Consequently, the local prototypes evolve dynamically during local model training. To submit more representative prototypes, each client regenerates them via the local model parameters $\textit{{w}}_{m,t}^{(E)}$. Formally, let $\textit{{c}}^{k}_{m,t}$ be the submitted local prototype, is computed as:
$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$
where $\textit{{r}}_{m,t}^{(E)}$ is the parameters of feature extractor.
Normalization
Considering that malicious clients amplify their submitted local prototypes, the local prototype of each class in each client is normalized. Formally, the local prototype is normalized by the following formula:
$\widetilde{\textit{{c}}}^{k}_{m,t}=\textit{{c}}^{k}_{m,t}/\|\textit{{c}}^{k}_{m,t}\|,\forall m\in\mathcal{S},k\in\mathcal{I},$
where $\widetilde{\textit{{c}}}^{k}_{m,t}$ is a unit vector. After normalization, each local prototype needs to be encrypted for privacy protection.
Input: $\textit{{r}}_{m,t}^{(E)}$, $\mathcal{D}_{m}$
Output: $\{\llbracket\textit{{c}}_{m,t}^{k}\rrbracket\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
1 for each class $k$in $\mathcal{I}$ do
2       $\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}^{|\mathcal{D}^{k}_{m}|}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)})$;
3       $\widetilde{\textit{{c}}}^{k}_{m}=\textit{{c}}^{k}_{m,t}/\|\textit{{c}}^{k}_{m,t}\|$;
4       Encrypt $\widetilde{\textit{{c}}}^{k}_{m,t}$ by Verifier’s $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$;
5
return $\{\llbracket\textit{{c}}_{m,t}^{k}\rrbracket\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
Algorithm 4 Handling
Encryption
To protect privacy of clients, they encrypts their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
5.3 Secure Aggregation
To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to aggregate global prototypes. The protocol consists of normalization verification and secure two-party computation.
5.3.1 Normalization verification
Since malicious participants may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized. Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product. Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1. If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$. After validation, Verifier sends the set $\mathcal{S}$ to Aggregator.
5.3.2 Secure Two-party Computation
To achieve privacy preservation and Byzantine-robust aggregation results, two servers perform secure two-party computation to calculate global prototypes. The computing of global prototype is shown in Algorithm 5 and described as follows.
Input: $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, $\mathcal{S}$, $\chi$.
Output: $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$.
1 Aggregator:
2      Compute trusted prototype $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ by formula (5);
3       Compute $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ by formula (6);
4       Compute $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket_{Pkv}$ and $\llbracket\chi^{\prime}\rrbracket_{Pkv}$;
5       $\llbracket h^{k}_{m,t}\rrbracket_{Pkv}\leftarrow\textbf{OutPut}(\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket_{Pkv},\llbracket\chi^{\prime}\rrbracket_{Pkv})$;
6       Randomly select a $n$-dimensional vector $\textit{{V}}^{n}$ and a number $p$;
7       Send $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ to Verifier;
8
9Verifier:
10      Decrypt $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ with $Skv$;
11       Compute $j^{k}_{m,t}$ by formula (8);
12       Compute $\textit{Sum}^{k}_{t}$;
[…]
14       Send $\textit{Sum}^{k}_{t}$, $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$, and $\llbracket\textit{{V}}^{n}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to Aggregator;
15
16 Aggregator:
17      Compute $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}\leftarrow\frac{1}{\textit{{V}}^{n}}\odot\llbracket\textit{{V}}^{n}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$;
18       Aggregate global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by formula (9);
19       Distribute global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ to clients;
Algorithm 5 SecComput
Specifically, malicious clients exploit poisoned local data to generate local prototypes with low credibility. Conversely, local prototypes submitted by benign clients should have high credibility. However, for each class without a trusted prototype direction, it is difficult to assess the credibility of submitted prototypes. To this end, Aggregator computes the trusted prototype via the formula:
$\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\llbracket\textit{{c}}_{m,t}^{k}\rrbracket_{Pkv},\forall k\in\mathcal{I},$ (5)
where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round. $\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: 1) In terms of magnitude, this is because the normalized malicious prototype do not affect the magnitude of trusted prototype, only the direction of trusted prototype; 2) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients. Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation. To achieve this, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result. Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype. If the direction of local prototype is similar to that of trusted prototype, its credibility is higher. Formally, Aggregator computes the credibility of local prototype by the following formula,
$\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ $=\llbracket\textit{sim}(\widetilde{\textit{{c}}}^{k}_{m,t},\textit{{C}}^{\prime k}_{t+1})\rrbracket_{Pkv}=\left\llbracket\frac{\widetilde{\textit{{c}}}^{k}_{m,t}\cdot\textit{{C}}^{\prime k}_{t+1}}{\|\widetilde{\textit{{c}}}^{k}_{m,t}\|\|\textit{{C}}^{\prime k}_{t+1}\|}\right\rrbracket_{Pkv}$ (6)
$=\frac{1}{\|\textit{{C}}^{\prime k}_{t+1}\|}\left(\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\right),$
where $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ denotes the credibility of local prototype. Additionally, we set the detection threshold $\chi$ and define local prototypes with its credibility less than $\chi$ as anomalous local prototypes. During global prototype aggregation, we set the aggregation weight of anomalous local prototype to 0. Formally, the aggregation weight of each local prototype is calculated as follows:
$j^{k}_{m,t}=\begin{cases}0.&\textit{sim}^{k}_{m,t}<\chi\\ \textit{sim}^{k}_{m,t}.&\textit{sim}^{k}_{m,t}>\chi,\end{cases}$ (7)
where $j^{k}_{m,t}$ denotes the aggregation weight of local prototype for class $k$ at client $m$ in the $t$-th communication round. However, Aggregator cannot directly compare $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ with $\chi$ under encryption domain.
To implement comparison within encryption domain, the work in [40] offers a ciphertext comparison method that outputs the maximum value of homomorphic ciphertexts corresponding to two plaintexts in the range (0, 1) without decryption. However, since the cosine similarity ranges from -1 to 1, and the method cannot be applied directly. To resolve the range mismatch, Aggregator sets $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket=\frac{1}{2}(\llbracket\textit{sim}^{k}_{m,t}\rrbracket+\llbracket 1\rrbracket)$ and $\chi^{\prime}=\frac{1}{2}(\chi+1)$ as input to the comparison method [40]. This is because the size relationship between $\llbracket\textit{sim}^{k}_{m,t}\rrbracket$ and $\chi$ remains unchanged. Then, Aggregator encrypts $\chi^{\prime}$ to get $\llbracket\chi^{\prime}\rrbracket$, and calculates the maximum value $\llbracket h^{k}_{m,t}\rrbracket$ between $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $\llbracket\chi^{\prime}\rrbracket$ by Algorithm 6.
Input: $(\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,\llbracket\chi^{\prime}\rrbracket)\in(\llbracket 0\rrbracket,\llbracket 1\rrbracket),d\in\mathbb{N}$
Output: an max value of $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ or $\llbracket\chi^{\prime}\rrbracket$
1 $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,b=\llbracket\chi^{\prime}\rrbracket$;
2 $q_{1}=\frac{(a+b)}{2},q_{2}=\frac{(a-b)}{2}$;
3 $a_{0}=q_{2}^{2},b_{0}=q_{2}^{2}-1$;
4 for each $n$in $(0,d-1)$ do
5       $a_{n+1}=a_{n}(1-\frac{b_{n}}{2})$;
6       $b_{n+1}=b_{n}^{2}(\frac{b_{n}-3}{4})$;
7
[…]
return $(q_{1}+q_{3})$
Algorithm 6 OutPut
After obtaining the maximum value $\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$, Aggregator collaborates with Verifier to calculate the aggregation results. Specifically, Aggregator selects a random value $p$, and multiplies it by the maximum value to get $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$. Additionally, Aggregator chooses a random vector V with the same dimension as the prototype, and computes its hadamard product with local prototype to obtain $\textit{{V}}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, where $\odot$ denotes the hadamard product. Then, Aggregator sends $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$ and $\textit{{V}}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ to Verifier. The Verifier then decrypts these values by its secret key $Skv$ to get $p\times h^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$. Since the random value $p$ and random vector V obfuscate $h^{k}_{m,t}$ and $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, respectively, Verifier cannot extract sensitive information about clients from $p\times h^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$. Then, the aggregation weight is rewritten as the following formula:
$j^{k}_{m,t}=\begin{cases}0.&\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}\\ p\times h^{k}_{m,t}&\textit{Round}(p\times h^{k}_{m,t},6)>\textit{min}^{k}_{t},\end{cases}$ (8)
where $\textit{Round}(p\times h^{k}_{m,t},6)$ denotes that $p\times h^{k}_{m,t}$ is rounded to the 6-th decimal place. This is because CKKS decrypts the ciphertext with an error in the range of $10^{-7}$ [41]. Moreover, $\textit{min}^{k}_{t}$ denotes the smallest value for class $k$ in set $\{\textit{Round}(p\times h^{k}_{m,t},6)\}_{m\in\mathcal{S}}^{|\mathcal{S}|}$. The $\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}$ means that the local prototype $\textit{{c}}^{k}_{m,t}$ satisfies $\textit{sim}^{k}_{m,t}<\chi$, thus its aggregation weight $j^{k}_{m,t}$ is 0. Additionally, Verifier calculates $\textit{Sum}^{k}_{t}=\sum_{m\in\mathcal{S}}j^{k}_{m,t}$ and encrypts $j^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using the client’s public key $Pkx$ to get $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$ and $\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$, and sends them to Aggregator.
The Aggregator computes $\frac{1}{\textit{{V}}}\odot\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$, and aggregates the encrypted local prototypes to get the encrypted global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by the following formula,
$\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ $=\frac{1}{\textit{Sum}^{k}_{t}}\sum_{m\in\mathcal{S}}\llbracket j^{k}_{m,t}\rrbracket_{Pkx}\times\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}.$ (9)
Subsequently, the Aggregator distributes the encrypted global prototype to Clients. After the above process is completed, FL executes the next communication round until a predefined number of rounds is reached.
6 Analysis
In the section, we provide both convergence analysis and privacy analysis for PPFPL. Specifically, we make the following assumptions similar to existing general frameworks [23][42] for loss function (3).
Assumption 1.
Each loss function is $L_{1}$ Lipschitz smooth, which means that the gradient of loss function is $L_{1}$ Lipschitz continuous, we can get
$\begin{gathered}\left\|\nabla\mathcal{L}_{m,t}^{(e_{1})}-\nabla\mathcal{L}_{m,t}^{(e_{2})}\right\|_{2}\leq L_{1}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2},\end{gathered}$
where $\mathcal{L}_{m,t}^{(e_{1})}$ denotes loss function at the $(tE+e_{1})$-th local iteration in client $m$. This implies the following quadratic bound,
$\mathcal{L}_{m,t}^{(e_{1})}$ $-\mathcal{L}_{m,t}^{(e_{2})}$
$\leq\left\langle\nabla\mathcal{L}_{m,t}^{(e_{2})},\left(\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2}^{2}.$
Assumption 2.
The stochastic gradient $\textit{{g}}_{m,t}^{(e)}=\nabla\mathcal{L}\left(\textit{{w}}_{m,t}^{(e-1)};\mathcal{D}_{m}^{(e)}\right)$ is an unbiased estimator of the local gradient for each client. Suppose its expectation
$\mathbb{E}_{\mathcal{D}_{m}^{(e)}\sim\mathcal{D}_{m}}\left[\textit{{g}}_{m,t}^{(e)}\right]=\nabla\mathcal{L}\left(\textit{{w}}_{m,t}^{(e)};\mathcal{D}_{m}^{(e)}\right)=\nabla\mathcal{L}_{m,t}^{(e)},$
and its variance is bounded by $\sigma^{2}$:
$\mathbb{E}[\|\textit{{g}}_{m,t}^{(e)}-\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e)})\|_{2}^{2}]\leq\sigma^{2}.$
Based on the above assumptions, we have the following theorem and corollaries. Notably, we add “$\frac{1}{2}$” into the local iteration, denoted as $\{\frac{1}{2},1,\cdots,E\}$ in our analysis. For example, $tE$ denotes the time step before local prototype aggregation, and $tE+\frac{1}{2}$ denotes the time step between local prototype aggregation and the first local iteration of the $t$-th round.
Theorem 1.
In PPFPL, regardless of the percentage of malicious clients, for the $t$-th communication round, the variation of loss function for each benign client can be bounded as,
$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
where $G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
Proof.
Assuming that Assumption 1 holds, we can get
$\mathcal{L}_{m,t}^{(1)}$ $\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}+\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\left(\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}$ (10)
[…]
where (a) follows from $\textit{{w}}_{m,t}^{(1)}=\textit{{w}}_{m,t}^{(\frac{1}{2})}-\eta\textit{{g}}_{m,t}^{(\frac{1}{2})}$. Taking expectation on both sides of formula (LABEL:102218311), we can get
$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-\eta\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\rangle\right]+\frac{\eta^{2}L_{1}}{2}\mathbb{E}\left[\left\|\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}\right]$
[…]
$=\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\|\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})}\|_{2}^{2}+\frac{\eta^{2}L_{1}}{2}Var(\textit{{g}}_{m,t}^{(\frac{1}{2})})$
[…]
where (b), (d) and (e) follow from Assumption 2, (c) follows from $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$. Then, during the local computation step, the loss function is updated $E$ times, and it can be bounded as:
$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}.\$ (11)
Additionally, since a single communication round involves both local computation and secure aggregation, we need to compute the impact of the aggregation result for loss function of each benign client. Specifically, the loss function of each benign client at the $((t+1)E+\frac{1}{2})$ time step is represented as follows:
$\mathcal{L}_{m,t+1}^{\frac{1}{2}}=\mathcal{L}_{m,t}^{E}+\mathcal{L}_{m,t+1}^{\frac{1}{2}}-\mathcal{L}_{m,t}^{E}$ (12)
$\stackrel{{\scriptstyle}}{{=}}\mathcal{L}_{m,t}^{E}+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+2})-\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+1})$
$\overset{}{=}\mathcal{L}_{m,t}^{E}-\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+2})+\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+1})$
[…]
where (f) follows from $-1\leq\textit{sim}(\cdot,\cdot)\leq 1$. Taking expectation on both sides of formula (12), we can get
$\mathbb{E}\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]$ $\leq\mathcal{L}_{m,t}^{E}+2\lambda.$ (13)
Thus, during the $t$-th communication round, according to the formula (11) and formula (13), the variation of loss function for each benign client can be bounded as,
$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
where
$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
Thus, Theorem 1 is proved. ∎
Corollary 1.
Given any fixed $\lambda$ and $E$, the $G(\eta)$ is a convex function with respect to $\eta$.
Proof.
To prove that $G(\eta)$ is a convex function for $\eta$, we need to prove that the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative. Thus, we have
$\frac{dG(\eta)}{d\eta}=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2},$
and
$\frac{d^{2}G(\eta)}{d\eta^{2}}=L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\sigma^{2}.$
Since $L_{1}$, $E$, and $\sigma^{2}$ are all greater than 0, we have $\frac{d^{2}G(\eta)}{d\eta^{2}}\textgreater 0.$ Thus, $G(\eta)$ is proved to be a convex function and there exists a minimum value of $G(\eta)$. ∎
Corollary 2.
Given any fixed $\lambda$ and $E$, the variation of the loss function compared to the previous round exists a minimum bound when $\eta=\eta^{\star}$, where $\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}$.
Proof.
When $G^{\prime}(\eta)$ equals 0, then $G(\eta)$ obtains a extremum value. Let $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=0,$ so we get
$(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2}=0.$
Thus, $G(\eta)$ exists a extremum value when $\eta=\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}.$ Because corollary 1 proves the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative, the extremum value is the minimal value. Therefore, we can understand that the variation of loss function exists a minimum bound when $\eta=\eta^{\star}$. ∎
Corollary 3.
Given any fixed $\eta$ and $E$, the loss function of arbitrary client monotonously decreases in each communication round when $\lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.$
Proof.
To guarantee that the local loss function decreases after each communication, we need to make sure that $G(\lambda)\textless 0$, so we have
$-\left(\eta-\frac{L_{1}\eta^{2}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda\textless 0.$
After simplification, we get
$\lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.$
Theorem 2.
The Aggregator, Verifier, and malicious clients cannot access any sensitive information about benign clients.
Proof.
During the secure aggregation, the two servers can obtain plaintext information $\|\textit{{C}}^{\prime k}_{t+1}\|$ and $\textit{Sum}^{k}_{t}$, where $\|\textit{{C}}^{\prime k}_{t+1}\|$ denotes the norm of trusted update, $\textit{Sum}^{k}_{t}$ denotes the sum of aggregation weight of class $k$. For non-colluding Aggregator and Verifier, they cannot get any sensitive information from these plaintext information. In addition, for colluding malicious clients, when there are $(|\mathcal{S}|-1)$ malicious clients in PPFPL, they can infer the local prototype about benign client from the encrypted global prototype. However, the real scenario does not exist when there are $(|\mathcal{S}|-1)$ malicious clients. Since benign clients receive only global prototypes distributed by Aggregator, they cannot deduce information about others. Therefore, neither third-party entities nor malicious clients can derive sensitive information about benign clients. ∎
7 Experiments
In this section, we evaluate the performance of PPFPL in the presence of data poisoning attacks on Non-IID data.
7.1 Experimental Settings
7.1.1 Datasets and Models
Similar to previous works [18][16][43], we utilize three public available datasets, namely MNIST, FMNIST, and CIFAR10, to evaluate performance of our PPFPL.
  1. 1.
    MNIST: This dataset comprises handwritten digits, with a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each represented as a 28 $\times$ 28 gray-scale image.
  2. 2.
    FMNIST: The FMNIST is a clothing image dataset containing a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each of which is a 28$\times$28 grey-scale image.
  3. 3.
    CIFAR10: As a color image dataset, CIFAR10 contains a training set of 50,000 samples in 10 classes and a test set of 10,000 samples, with each image measuring 32$\times$32 pixels.
Furthermore, we apply CNN as local model to both MNIST and FMNIST. ResNet18 is used for CIFAR10, where we initialize the local model with pre-trained parameters. These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
7.1.2 Settings of FL
We employ a cross-silo configuration in FL setup. Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round. The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively. We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to 5 by default. These hyperparameters are consistent across all clients.
7.1.3 Non-IID Settings
To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios. Specifically, large organizations (e.g., hospitals, companies) possess different data classes, and their class distributions may differ significantly, or even be missing some classes altogether. When these organizations participate in federated training, the union of their data classes defines the entire FL classification task. This phenomenon leads to Non-IID data across organizations. The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a certain percentage of all classes. However, in cross-silo scenarios, this contradicts the assumptions of Dirichlet distribution.
To model the data distribution in cross-silo scenarios, we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts. In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity. Clients are randomly assigned classes, with partial class overlap among them. To visualize the different data distributions, we plot heat maps as shown in Fig. 2.
((a)) Avg=3, Std=2
((b)) Avg=4, Std=2
((c)) Avg=5, Std=2
Fig. 2: Heat maps for different data distributions under heterogeneity distribution.
7.1.4 Setting of Data Poisoning Attacks
In our experiments, we consider two types of data poisoning attacks: feature attacks and label attacks. For feature attacks, malicious clients randomly alter their own training data features in a completely randomized manner without following any specific rule. For label attacks, malicious clients alter the labels of their own training data to incorrect labels. All training data of malicious clients are modified. Malicious clients do not dynamically change their behaviors during federated training, and the attack remains consistent throughout the training period. Since the number of malicious clients affects FL performance differently, we set different proportions of malicious clients. Formally, Att is set as the proportion of malicious clients.
7.1.5 Evaluation Measure
Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Therefore, we evaluate the FL performance by testing the average accuracy of benign clients. For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks. In addition, we compare our PPFPL with robust schemes (i.e., Krum [44] and Foolsgold [45]) and privacy-preserving robust schemes (i.e., ShieldFL [18] and PBFL [16]) under the same experiment conditions.
7.2 Experimental Results
7.2.1 Visualization of Prototypes
We use t-SNE to visualize client-submitted prototypes in PPFPL, as shown in Fig. 3. Specifically, we display prototypes from clients under feature and label attacks with Att =20% on the CIFAR10 dataset. This figure shows the distribution of these prototypes in a single communication round. We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks. In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area. This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas. To mitigate the severe impact of malicious clients, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results. The following is the security evaluation of PPFPL.
((a)) Feature attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
((b)) Feature attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
((c)) Label attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
((d)) Label attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
Fig. 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
7.2.2 Security Evaluation
((a)) Feature attacks, FMNIST.
((b)) Feature attacks, CIFAR10.
((c)) Label attacks, FMNIST.
((d)) Label attacks, CIFAR10.
Fig. 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
((a)) Feature attacks, CIFAR10.
((b)) Label attacks, CIFAR10.
Fig. 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$. Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$. Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, and PBFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
The experiment results are shown in Fig. 4, where the black dashed line represents the highest accuracy of PPFPL without attacks. We observe that the accuracies of ShieldFL and PBFL (under attacks) are severely lower than FedAvg (no attack), especially for CIFAR10. This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious. This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance. In contrast, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence. This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients. Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
Additionally, we evaluate the change of loss during training for PPFPL and the existing schemes, as shown in Fig. 5. We notice that the loss of PBFL fluctuates significantly during training process, while PPFPL is relatively smooth. This indicates that data poisoning attacks disrupt the convergence of PBFL, while the convergence of PPFPL is not affected by data poisoning attacks, thus ensuring the accuracy of the model. This is consistent with our Corollary 3. Specifically, PPFPL still satisfies convergence in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them. Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security. Notably, our experiments show that the performance impact of feature attacks or label attacks is similar for PPFPL and other schemes. This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
((a)) Feature attacks
((b)) Label attacks
Fig. 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
7.2.3 Different Data Distributions for PPFPL
To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10. In addition, we compare the performance of Krum, Foolsgold, ShieldFL, and PBFL. The specific experimental results are shown in Table 1. Notably, since test average accuracy has fluctuated after each communication round, we are select the average of five highest test average accuracies across the communication rounds. From Table 1, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution. This is attributed to that the client-submitted prototype does not change due to the change in data distribution. In addition, we observe a slight degradation in the performance of PPFPL with higher proportion of attacks. This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon. Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
Table 1: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
Dataset Method Att % Std Test Average Accuracy (%)
Avg = 3 Avg = 4 Avg = 5 Avg = 6 Avg = 7
MNIST Krum 20 1 95.96 96.95 97.01 97.04 97.12
2 95.67 94.28 94.88 95.51 96.20
30 1 95.74 96.61 96.82 96.78 96.92
2 95.67 94.28 94.88 94.77 95.80
Foolsgold 20 1 96.69 94.03 95.13 95.76 95.39
2 96.19 94.82 96.73 95.27 95.75
30 1 96.47 93.76 94.98 95.21 95.00
2 96.02 94.74 96.58 95.01 95.49
PBFL 20 1 97.78 97.57 96.34 97.10 96.98
2 97.36 96.17 96.26 94.64 95.41
30 1 95.78 94.15 95.95 96.50 96.75
2 95.47 93.62 93.48 94.70 94.62
ShieldFL 20 1 96.91 94.05 95.14 95.64 95.25
2 97.25 94.88 95.69 95.44 95.76
30 1 96.77 93.79 94.93 95.60 95.81
2 97.08 94.69 95.46 95.73 94.46
PPFPL 20 1 97.87 97.53 97.35 97.08 96.89
2 97.74 97.49 97.43 96.82 96.56
30 1 97.85 97.04 97.02 96.52 96.37
2 97.52 96.92 96.70 96.50 96.48
FMNIST Krum 20 1 84.24 84.80 86.03 86.52 87.35
2 85.02 83.64 86.82 84.75 86.08
30 1 83.72 84.38 85.49 85.12 86.79
2 84.62 82.73 85.73 82.86 84.09
Foolsgold 20 1 84.88 84.73 84.04 83.55 82.62
2 82.49 83.88 85.31 81.75 82.24
30 1 84.06 84.01 83.50 82.70 82.04
2 81.86 83.43 84.98 81.19 82.68
PBFL 20 1 87.01 80.37 76.93 74.26 72.10
2 87.42 76.31 74.47 72.99 72.43
30 1 86.92 75.45 73.06 73.70 72.84
2 79.34 72.68 72.81 70.57 70.49
ShieldFL 20 1 87.01 78.90 78.06 70.90 70.26
2 76.47 77.89 70.96 67.47 66.80
30 1 86.24 78.46 77.51 69.33 68.81
2 75.81 77.35 70.57 66.42 65.46
PPFPL 20 1 91.38 90.18 89.27 88.56 88.13
2 90.48 89.63 88.45 88.79 88.20
30 1 90.97 90.08 88.93 88.33 88.14
2 90.40 88.62 87.74 88.41 87.97
CIFAR10 Krum 20 1 62.91 61.24 68.19 68.87 69.22
2 57.90 57.94 68.33 68.65 68.79
30 1 60.10 60.73 66.36 68.70 69.01
2 56.67 57.29 67.03 68.49 68.51
Foolsgold 20 1 60.91 59.24 66.19 66.45 66.52
2 56.90 55.94 66.33 66.62 66.70
30 1 60.37 58.68 65.64 66.26 66.44
2 56.17 55.09 65.85 66.40 66.41
PBFL 20 1 38.06 39.18 41.81 42.42 42.75
2 39.46 40.54 41.30 42.12 42.76
30 1 37.18 39.02 41.39 42.32 42.41
2 37.80 37.52 41.08 41.96 42.67
ShieldFL 20 1 61.62 61.94 68.31 69.53 68.66
2 57.84 57.26 66.14 67.43 68.62
30 1 60.94 61.25 67.50 67.01 68.24
2 57.10 56.76 65.47 66.42 67.69
PPFPL 20 1 83.41 80.95 77.82 74.76 71.35
2 83.44 83.24 77.19 74.64 71.60
30 1 82.31 81.03 77.18 74.53 71.15
2 83.16 82.37 77.47 74.62 71.19
7.2.4 Secure Aggregation Protocol for PPFPL
To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). In addition, the $\boldsymbol{X}$ in Table 2 denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging. From Table 2, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings, but the magnitude of that degradation is limited. This is because each client minimizes its local classification loss and aligns local prototype with global prototype during local model training. However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss. Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients. Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which suggests that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol. The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks?
Table 2: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
      Datasets       Att       $\chi$ = -1       $\chi$ = 0       $\chi$ = 0.2       $\chi$ = 0.5       $\boldsymbol{X}$
      MNIST       10%       96.23       97.86       97.62       98.16       95.73
      20%       96.10       97.20       97.55       97.03       94.35
      30%       95.21       97.10       97.47       97.82       94.20
      40%       94.54       96.75       96.51       96.48       93.14
      FMNIST       10%       88.37       90.68       90.23       90.78       86.14
      20%       88.20       90.28       90.36       90.48       86.12
      30%       88.03       90.14       90.06       90.16       85.71
      40%       87.62       90.23       90.41       90.04       85.29
      CIFAR10       10%       81.66       83.59       83.10       83.24       80.97
      20%       81.61       83.57       83.67       83.31       79.53
      30%       81.38       83.45       83.35       83.21       78.08
      40%       80.74       83.05       83.27       82.94       78.00
((a)) Att = 20%
((b)) Att = 40%
((c)) Att = 60%
((d)) Att = 80%
Fig. 7: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
7.2.5 High Proportion Attacks for PPFPL
To test the performance of PPFPL under high proportion attacks, we evaluate the performance of PPFPL under Att = 20%, 40%, 60% and 80%, respectively. In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). The $\lambda$ denotes the importance weight of the auxiliary term in loss function, which can be considered as the degree of influence among clients. The experimental results are shown in Fig. 7. We surprisingly observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients. This is because the benign client does not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data. Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets. On the contrary, when Att is 60% or 80%, the accuracy tends to decrease with the increase of $\lambda$. This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients. Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but increases resistance to data poisoning attacks. Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
7.2.6 Stability of PPFPL
((a)) Feature attacks
((b)) Label attacks
Fig. 8: Stability of PPFPL under MNIST, FMNIST, and CIFAR10.
It is observed from Fig. 4 and 5 that the performance of PPFPL fluctuates. Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments. To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of participants being malicious. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). Each experiment is repeated 10 times. The utilization of boxplots in Fig. 8 provides a clear visualization of the stability of PPFPL. The results show that PPFPL’s performance fluctuates less and remains stable against data poisoning attacks. The stability of PPFPL proves its reliability in real-world scenarios.
7.2.7 Scalability of PPFPL
Federated learning is often deployed with an uncertain number of clients, and scalability directly determines whether PPFPL can maintain stable performance in environments with different numbers of clients. To evaluate the scalability of PPFPL, we conducted experiments on the MNIST, FMNIST, and CIFAR10 datasets with 20% malicious clients, and the number of clients set to 20, 25, 30, 35, and 40. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=1$). Fig. 9 shows the scalability of PPFPL against data poisoning attacks. It is observed that the performance of PPFPL drops by 0.24% on MNIST and 0.42% on CIFAR10 when the number of participants is increased by 100%, which is normal in federated learning. Thus, PPFPL shows excellent scalability against data poisoning attacks with data heterogeneity.
((a)) MNIST
((b)) FMNIST
((c)) CIFAR10
Fig. 9: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
7.2.8 Impact of Local Iterations
We investigate the effect of the number of local iterations on the test average accuracy for PPFPL. We conduct experiments on MNIST, FMNIST, and CIFAR10 datasets, where 20% of malicious clients launched feature attacks. The number of local iterations are 1, 5, 10, 15, 20 and 25. The results are shown in Fig. 10. We observe that when the number of local iterations is 1, the changes in the local model are very small, resulting in slow training and relatively low accuracy for the same number of communication rounds. When the number of local iterations is 5, PPFPL has the highest test average accuracy on the MNIST dataset. However, on CIFAR10, the optimal number of local iterations should be set to 15. In addition, when the number of local iterations is too high, the test average accuracy decreases, which is due to the overfitting of the local model.
((a)) MNIST.
((b)) FMNIST.
((c)) CIFAR10.
Fig. 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
7.2.9 Efficiency Evaluation
((a))
((b))
Fig. 11: Communication parameters. (a) The number of parameters submitted by clients under different local models. (b) The number of parameters submitted by client using CNN model in one round under different scheme.
To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients in PPFPL, as shown in Fig. 11. We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same parameters. Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture. This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, not the model architecture. Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
Table 3: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
       Operation        PBFL        ShieldFL        PPFPL
       Encrypt        1.72s        1.42s        0.62s
       Decrypt        1.35s        1.19s        0.40s
In addition, ciphertext operations are the most important factor affecting the efficiency of our framework. We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations. Note that PBFL is a federated learning scheme based on CKKS, while ShieldFL is built upon the two-door HE. For each client, we use the time spent on encryption and decryption for each iteration as a metric. As shown in Table 3, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL. This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
7.3 Complexity Analysis
Table 4: Computation Overload and Communication Overload of client.
Scheme Computation Overload Communication Overload
PEFL [15] $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})$ $\mathcal{O}(gP_{h})$
ShieldFL [18] $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})$ $\mathcal{O}(gP_{h})$
PBFL [16] $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})$ $\mathcal{O}(gP_{h})$
PPFPL $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ $\mathcal{O}(nP_{h})$
We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15]. The results are shown in Table 4. The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption. Formally, the computational overhead is expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ denotes the prototype parameters, $T_{ch}$ denotes the time overload of encryption, $T_{tr}$ denotes the time overhead of local model training, and $T_{pro}$ denotes the time overhead of prototype generation. Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$. Other similar schemes [16][18][15] adopt homomorphic encryption on gradients, their encryption overhead is $\mathcal{O}(gT_{ch})$, where $g$ denotes the gradient parameters. Since $g>p$, the overload satisfies $\mathcal{O}(gT_{ch})>\mathcal{O}(pT_{ch})$. Hence, the relationship $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ holds. Consequently, PPFPL imposes a reasonable computational overhead on clients.
In addition, the client communication overhead is $\mathcal{O}(nP_{h})$ in PPFPL, where $P_{h}$ denotes the communication complexity of a number. Compared with [16][18][15], the relationship $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds.
8 Concludes
In this paper, we propose PPFPL, a privacy-preserving federated prototype learning framework, which improves performance while resisting data poisoning attacks in poisoned Non-IID data. Our work is the first to introduce prototype learning into PPFL to address privacy threats and data poisoning attacks in across-silo FL. As for how to implement lightweight privacy-preserving FL to resist data poisoning attacks on Non-IID data, we leave them to future exploration.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grants 62272256 and 62202250, the Major Program of Shandong Provincial Natural Science Foundation for the Fundamental Research under Grant ZR2022ZD03, the National Science Foundation of Shandong Province under Grant ZR2021QF079, the Talent Cultivation Promotion Program of Computer Science and Technology in Qilu University of Technology (Shandong Academy of Sciences) under Grant 2023PY059, the Pilot Project for Integrated Innovation of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences) under Grant 2022XD001, and the Colleges and Universities 20 Terms Foundation of Jinan City under Grant 202228093.
CRediT authorship contribution statement
Hongliang Zhang: Writing – original draft, Validation, Software, Methodology, Conceptualization. Jiguo Yu: Writing – review & editing, Supervision. Fenghua Xu: Software, Validation, Writing – review & editing. Chunqiang Hu: Writing – review & editing. Yongzhao Zhang: Writing – review & editing, Supervision. Xiaofen Wang: Software, Validation. Zhongyuan Yu: Validation. Xiaosong Zhang: Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.

### v4

@@
 Abstract
-Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train a deep learning model by submitting hidden model updates.
-Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients.
-Existing solutions have struggled to improve the performance of cross-silo PPFL in poisoned Non-IID data.
-To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while effectively resisting data poisoning attacks.
-Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of tampered data distribution on federated learning.
-Moreover, we utilize two servers to achieve Byzantine-robust aggregation by secure aggregation protocol, which greatly reduces the impact of malicious clients.
-Theoretical analyses confirm the convergence of PPFPL, and experimental results on publicly available datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID conditions.
-keywords:
-Privacy-preserving, cross-silo, federated learning, data poisoning attacks, poisoned Non-IID data.
-journal: Nuclear Physics B
-1 Introduction
-Federated Learning (FL) is a distributed learning paradigm where each client submits its model updates instead of raw training data.
+Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train by submitting hidden model updates.
+Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients in cross-silo scenarios.
+Existing solutions have struggled to improve the performance of PPFL in poisoned Non-IID data.
+To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while protecting the privacy of clients.
+Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distribution.
+In addition, we design a secure aggregation protocol by utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, which greatly reduces the impact of malicious clients.
+Theoretical analyses confirm the convergence and privacy of PPFPL, and experimental results on publicly datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID data.
+Index Terms:
+Privacy-Preserving, Federated Learning, Cross-Silo, Data Poisoning Attacks, Poisoned Non-IID Data.
+I Introduction
+Federated Learning (FL) is a distributed learning paradigm where each client shares its model updates instead of raw training data.
 In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1].
 To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3].
-Specifically, most clients in cross-silo FL are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities.
-However, client-submitted model updates are vulnerable to privacy inference attacks, which threats the privacy security of cross-silo FL [4][5].
-To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] or Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed.
-In particular, the DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the accuracy of the global model by introducing noise.
-Conversely, the HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
-While HE-based works [8] [9] demonstrated their effectiveness in privacy preservation within cross-silo FL, they are susceptible to data poisoning attacks due to the distributed training paradigm [10][11].
-Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, severely impacting the performance of PPFL [12][13].
-Moreover, in PPFL, data poisoning attacks are more concealed due to privacy technology obscuring model updates from malicious clients, which raises the difficulty of defense.
-To audit obscured model updates, recent works [14][15][16][17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious updates in ciphertext.
-However, these works overlook the heterogeneity of local training data (i.e., Non-IID) issue in FL, which causes distinct inconsistency of benign updates, thereby disrupting the defense against malicious updates.
-To distinguish between benign and malicious updates in the Non-IID data, existing solutions [18][19] incorporate clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase.
-While these solutions are effective against model poisoning attacks, they are inadequate for countering data poisoning attacks.
-This limitation stems from their reliance on specialized training techniques designed for Non-IID data (e.g., FedProx[20], FedDyn[21], or FedLC[22], etc.), which enhance the performance of global model in Non-IID scenarios.
-However, in data poisoning attacks, malicious clients deliberately tamper with the features and labels of their training data, thereby forming poisoned Non-IID data.
-These manipulations distort the optimization direction of local model training and significantly degrade the effectiveness of the specialized training techniques.
-Consequently, integrating the above defense solutions [18][19] with such training techniques fails to effectively mitigate data poisoning attacks in poisoned Non-IID data.
-Therefore, a critical challenge remains: how to improve the performance of cross-silo PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
-Inspired by prototype learning, several works [23][24] designed prototype aggregation to address the Non-IID issue in federated learning, where prototypes are transmitted between the server and clients.
-Each prototype represents a class feature that is calculated from the mean representations transformed from the samples belonging to the same class.
-For instance, the works in [23][24] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”.
-By exchanging these prototypes, clients can gain more knowledge about the concept of “cat”.
+Specifically, in cross-silo FL, the clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities.
+However, client-submitted model updates are vulnerable to privacy attacks, which threats the privacy security of cross-silo FL [4][5].
+To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] and Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed.
+In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by introducing noise.
+Conversely, HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
+While HE-based approaches demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11].
+Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13].
+Moreover, Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and more difficult to defend.
+To audit obscured model updates, existing methods [14, 15, 16, 17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates in ciphertext.
+However, these works overlook the data heterogeneity in FL, i.e., the data among clients is typically non-independent and identically distributed (Non-IID).
+The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17] to distinguish whether the deviations are caused by data heterogeneity or by malicious model updates.
+To distinguish between benign and malicious updates in the Non-IID data, existing methods [18, 19, 20] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase, thereby filtering out updates caused by model poisoning attacks.
+However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, thereby forming poisoned Non-IID data.
+These manipulations distort the optimization direction of local training, causing these methods [18, 19, 20] to optimize toward the tampered Non-IID distribution, thereby degrading the performance of FL.
+Therefore, a critical challenge remains: how to enhance the performance of PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
+Inspired by prototype learning [21], several works [22, 23, 24, 25] proposed prototype aggregation to address the Non-IID issue in FL, where prototypes are exchanged between the server and clients.
+Each prototype represents a class-level feature, computed as the mean of the feature representations of samples belonging to the same class.
+For instance, the work in [23] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”.
+By exchanging these prototypes, clients gain more knowledge about the concept of “cat”.
 Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
-In cases of poisoned Non-IID data, tampered data features and data distribution cannot be repaired since the server cannot manipulate the behavior of malicious clients.
-This prompts us to pose a question: Is it possible to design a PPFL that leverages prototype learning, so that client-submitted model updates are affected only by tampered data features and not by tampered data distribution, while incorporating a secure aggregation protocol to achieve Byzantine-robust results?
-To answer this question, we propose Privacy-Preserving Federated Prototype Learning framework, named PPFPL, which is suitable for cross-silo scenarios.
-Our core idea is to utilize prototype learning to address the issue of tampered data distribution, while designing a secure aggregation protocol across two servers to achieve Byzantine-robust aggregation results without compromising privacy.
-PPFPL employs two “non-colluding” servers who are “honest but curious” to execute a secure aggregation protocol that utilizes cosine similarity as a method to penalize malicious clients.
-In addition, HE and SMC are adopted as underlying technologies to add the privacy-preserving mechanism of the secure aggregation protocol.
-The main contributions of the paper are summarized as follows.
-  • 1.
-    We are the first to introduce prototype learning into PPFL to address data poisoning attacks in across-silo FL, which eliminates the impact of tampered data distribution for client-submitted model updates while significantly improving the FL performance in poisoned Non-IID data.
-  • 2.
-    We design a secure aggregation protocol across two servers to aggregate client-submitted prototypes with HE and SMC techniques, which ensures that two servers and malicious clients cannot utilize client-submitted prototypes to invade the privacy of benign clients.
-  • 3.
+[…]
+Our main contributions are summarized as follows.
+  • We are the first to introduce prototype learning into PPFL to defend against data poisoning attacks.
+By transferring prototypes between clients and servers, our framework mitigates the impact of tampered data distribution and enhances the FL performance in poisoned Non-IID data.
+  • We utilize HE and SMC techniques to design a secure aggregation protocol across two servers to filter prototypes submitted by malicious clients, ensuring Byzantine-robust results while protecting the privacy of benign clients.
   • We provide a convergence guarantee for PPFPL in the presence of data poisoning attacks, which theoretically ensures the framework’s feasibility.
-  • 4.
   • Compared to existing methods, the superiority of our framework has been empirically validated in poisoned Non-IID data.
 The rest of the paper is organized as follows.
-Section 2 reviews recent PPFL works against poisoning attacks.
-Section 3 introduces prototype learning and HE.
-Section 4 formalizes the system model of PPFPL.
-Our framework is presented in Section 5.
-Section 6 provide theoretical analysis.
-Section 7 reviews experimental results.
-Finally, Section 8 concludes this paper.
-2 RELATED WORK
-Privacy preservation in FL: Although across-silo FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of clients.
-To resist privacy attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates.
-Specifically, DP-based schemes [6][7][25][26] deploy local differential privacy into model updates to ensure privacy without compromising the utility of model updates.
-Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which reduces the performance of their global models to some extent.
-In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without degrading the global model’s accuracy.
+Section II reviews PPFL works against poisoning attacks.
+Section III introduces prototype learning and HE.
+Section IV formalizes the system model.
+Our PPFPL is detailed in Section V.
+Section VI provide theoretical analysis.
+Section VII reviews experimental results.
+Finally, Section VIII concludes this paper.
+II RELATED WORK
+Privacy preservation in FL: Although FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients.
+To resist such attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates.
+Specifically, DP-based schemes [26, 27, 28, 29] deploy local differential privacy into model updates, ensuring privacy without compromising the utility of model updates.
+Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which degrades the FL performance to some extent.
+In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy.
 Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [30].
-They proposed a multi-party machine learning scheme using Paillier [27] technique without compromising participants’ private data.
-Considering the heavy communication overhead of Paillier, the work in [9] proposed a privacy-preserving federated learning scheme using CKKS (i.e., Cheon-Kim-Kim-Song).
-This scheme significantly reduces computational overhead associated with ciphertexts.
+They proposed a multi-party machine learning scheme using Paillier [31] technique without compromising clients’ privacy.
+Considering the heavy communication overhead of Paillier, the work in [32] proposed a privacy-preserving FL using CKKS (i.e., Cheon-Kim-Kim-Song) that reduces computational overhead associated with ciphertexts.
 This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [33].
@@
 demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [34].
-Additionally, they proposed an aggregated defense strategy that can identify malicious clients in FL to circumvent data poisoning attacks.
-In addition, Zhang et al.
-proposed an effective framework capable of tracking the attack time, objective, type, and poisoned location for model updates [30].
+Additionally, they proposed an aggregated defense strategy that identifies malicious clients in FL to circumvent data poisoning attacks.
+Similarly, the works in [35, 36, 37] proposed detection methods to filter malicious model updates.
 Differently, Doku et al.
-employed SVM to audit client’s local training data for excluding malicious clients, which prevents tampered data from poisoning the global model [31].
-However, this violates the privacy of clients to some extent.
-Furthermore, in practice, the threats of privacy and data poisoning attacks usually coexist in across-silo FL.
-PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while also concealing data poisoning attacks from malicious clients.
-Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15] [14][16][17] combine HE and SMC to identify anomalous model updates in ciphertexts.
+employed SVM to audit client’s local training data for excluding malicious clients [38].
+However, this work violates the privacy of clients to some extent.
+Furthermore, the threats of privacy and data poisoning attacks usually coexist in practice across-silo FL.
+PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while concealing data poisoning attacks from malicious clients.
+Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17] combine HE and SMC to identify anomalous model updates in ciphertexts.
 For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function.
 Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks.
-However, these schemes [15] [14][16][17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious updates or Non-IID data.
+However, these schemes [15, 14, 16, 17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious model updates or Non-IID data.
 To audit malicious updates in PPFL with Non-IID data, the works in [18][19] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage.
@@
 They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[39], FedDyn[40], or FedLC[41], etc.) to radically improve performance of FL.
-These specialized techniques adds an auxiliary term into local loss function to constrain model updates of clients, which helps to increase consistency of model updates during local model training.
+These specialized techniques adds an auxiliary term into local optimization function to constrain model updates of clients, which helps to increase consistency of model updates during local model training.
 However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs.
-Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while improving the performance of the global model under poisoned Non-IID data.
+Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
 Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue.
-Specifically, the works in [23][24][32] are one of the first to propose federated prototype learning using the concept of prototype learning.
-Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), its core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes.
+Specifically, the works in [22, 23, 24, 25] are one of the first to propose federated prototype learning using the concept of prototype learning.
+Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes.
 In other words, each class holds its corresponding prototype that is independent of other classes.
-As a result, client-submitted prototype is affected by the samples and is independent of data distribution among clients.
+As a result, client-submitted prototype is affected by the samples but is independent of data distribution among clients.
 This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
 Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks.
-Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation in across-silo FL caused by tampered Non-IID data, while preserving the privacy of client-submitted prototypes.
-3 PRELIMINARIES
+Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation caused by tampered Non-IID data in across-silo FL, while preserving the privacy of client-submitted prototypes.
+III PRELIMINARIES
 This section introduces prototypes in federated learning and CKKS technology.
-3.1 Prototypes Meet Federated Learning
+III-A Prototypes Meet Federated Learning
 In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [42].
 Due to this property, the prototypes of the same class among clients are similar in FL task.
-Consequently, many FL schemes [23][32] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
+Consequently, many FL schemes [22, 23, 24, 25] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
 To further understand the prototype calculation in federated prototype learning, we introduce some basic notations below.
@@
 Subsequently, the server distributes the latest global prototypes to each client to further train their local model.
-3.2 CKKS
+III-B CKKS
 CKKS is a homomorphic encryption technology.
@@
 For a more detailed implementation principle and process of CKKS, please refer to paper [46].
-4 Problem Statement
+IV Problem Statement
 In the section, we formalize the PPFPL framework, define potential threats, and design goals.
-4.1 PPFPL framework
+IV-A PPFPL framework
 The framework of PPFPL consists of four entities, each of which has its specific function, as shown in Fig.
@@
 The specific functions of each entity are outlined as follows.
-Fig.
-1: The PPFPL framework.
+Figure 1: The PPFPL framework.
 ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients.
@@
 ④ Aggregator distributes encrypted global prototypes to Clients.
-  • 1.
   • Key Generation Center (KGC).
 The entity is responsible for generating and managing keys of both Clients and Verifier, which are important elements to ensure security of encryption/decryption process.
-  • 2.
   • Clients.
@@
 They has a pair of public/secret keys generated by KGC, denoted as $Pkx$/$Skx$.
-  • 3.
   • Aggregator. The Aggregator is a central server responsible for aggregating local prototypes submitted by clients.
-  • 4.
   • Verifier.
@@
 In our framework, we not only define the function of each entity, but also define potential threats.
-4.2 Potential Threats for PPFPL
+IV-B Potential Threats for PPFPL
 We discuss potential threats of PPFPL in detail.
@@
 In federated prototype learning, we define the following two representative types of data poisoning attacks from the perspective of data features and labels.
-  • 1.
   • Feature attacks.
@@
 Specifically, malicious clients tamper with the features of their training data to generate malicious prototypes for uploading, thereby affecting the training results of other clients.
-  • 2.
   • Label attacks.
 The malicious clients tamper with the labels of their training samples into other random labels to form the tampered data distribution, aiming to reduce the performance of the entire federated learning.
-4.3 Design Goals of PPFPL
+IV-C Design Goals of PPFPL
 The goal of our study is to improve performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks.
 Specifically, we design the PPFPL framework to fulfill the following goals:
-  • 1.
   • Security.
@@
 In other words, the model performance of each benign client is not affected by data poisoning attacks under different data distributions.
-  • 2.
   • Privacy. PPFPL should ensure privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
-  • 3.
   • Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar frameworks, thereby reducing privacy computation overheads and communication overheads.
-5 Design of PPFPL
+V Design of PPFPL
 In this section, we first provide the overview of PPFPL, and then describe each step in detail.
-5.1 Overview of PPFPL
+V-A Overview of PPFPL
 The execution process of PPFPL is summarized in Algorithm 1.
@@
 Specifically, PPFPL iteratively performs the following two steps:
-  • 1.
-    Step I.
-Local Computation: Each client $m$ trains its local model with its local dataset $\mathcal{D}_{m}$.
-Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
-  • 2.
+  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
   • Step II.
@@
 Algorithm 1 Overview of PPFPL
-5.2 Local Computation
-The local computation step includes two essential stages: local model training and prototype handling. The detail is outlined in Algorithm 2.
+V-B Local Computation
+The local computation step includes two essential stages: local model training and prototype handling stages. The detail is outlined in Algorithm 2.
 Input: $\mathcal{S}$, $\mathcal{D}_{m}$, $\eta$ , $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$, $E$.
@@
 Algorithm 2 Local Computation
-5.2.1 Local Model Training
+V-B1 Local Model Training
 During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype.
-To achieve this, an auxiliary term is incorporated into the local loss function.
-Formally, the local loss function of client $m$ is defined as:
+To achieve this, we design an auxiliary term by leveraging prototype learning in the local optimization function.
+Formally, the local optimization function of client $m$ is defined as:
 $\mathcal{L}(\textit{{w}}_{m};\mathcal{D}$ ,${}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})$ (3)
@@
 The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions.
-The objective of each client aims to minimize classification loss while keeping its local prototype close to the global prototype in the direction.
+The optimization function ensures that each client reduces the classification loss while aligning its local prototype with the global prototype in direction.
 Next, we present the process of local model training.
@@
 The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$.
-Firstly, each client $m$ uses the local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round.
+Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round.
 Subsequently, each client iteratively performs the following stages.
-  • 1.
   • In $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
-  • 2.
   • The client inputs the training data $\mathcal{D}_{m}^{(e)}$ into the local model’s feature extractor to compute the local prototype $\textit{{c}}^{k,(e)}_{m,t}$ via formula (1).
-  • 3.
   • The client computes the unbiased stochastic gradient by
@@
 Algorithm 3 Training
-5.2.2 Prototype Handling
+V-B2 Prototype Handling
 The prototype handling stage consists of three key phases: prototype generation, normalization, and encryption, as illustrated in Algorithm 4.
@@
 To protect privacy of clients, they encrypts their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
-5.3 Secure Aggregation
-To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to aggregate global prototypes.
+V-C Secure Aggregation
+To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to filter malicious local prototypes and obtain Byzantine-robust global prototypes.
 The protocol consists of normalization verification and secure two-party computation.
-5.3.1 Normalization verification
-Since malicious participants may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized.
+V-C1 Normalization verification
+Since malicious clients may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized.
 Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product.
@@
 If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$.
-After validation, Verifier sends the set $\mathcal{S}$ to Aggregator.
-5.3.2 Secure Two-party Computation
+After validation, Verifier sends the remaining client set $\mathcal{S}$ to Aggregator.
+V-C2 Secure Two-party Computation
 To achieve privacy preservation and Byzantine-robust aggregation results, two servers perform secure two-party computation to calculate global prototypes.
@@
 Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation.
-To achieve this, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result.
+To achieve the plaintext, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result.
 Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype.
@@
 After the above process is completed, FL executes the next communication round until a predefined number of rounds is reached.
-6 Analysis
+VI Analysis
 In the section, we provide both convergence analysis and privacy analysis for PPFPL.
@@
 $\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
-where $G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
+where
+$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
 Proof.
@@
 ∎
-7 Experiments
+VII Experiments
 In this section, we evaluate the performance of PPFPL in the presence of data poisoning attacks on Non-IID data.
-7.1 Experimental Settings
-7.1.1 Datasets and Models
+VII-A Experimental Settings
+VII-A1 Datasets and Models
 Similar to previous works [16][36][52], we utilize three public available datasets, namely MNIST, FMNIST, and CIFAR10, to evaluate performance of our PPFPL.
@@
     CIFAR10: As a color image dataset, CIFAR10 contains a training set of 50,000 samples in 10 classes and a test set of 10,000 samples, with each image measuring 32$\times$32 pixels.
-Furthermore, we apply CNN as local model to both MNIST and FMNIST.
-ResNet18 is used for CIFAR10, where we initialize the local model with pre-trained parameters.
+Furthermore, we apply Convolution Neural Network (CNN) as local model to both MNIST and FMNIST.
+For CIFAR-10, we employ ResNet18 as the local model, initialized with pre-trained parameters.
 These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
-7.1.2 Settings of FL
+VII-A2 Hyper-parameters Settings of FL
 We employ a cross-silo configuration in FL setup.
@@
 The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively.
-We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to 5 by default.
-These hyperparameters are consistent across all clients.
-7.1.3 Non-IID Settings
+We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to $E=5$.
+Except for special declared, the above default settings are applied.
+These hyper-parameters are consistent across all clients.
+VII-A3 Non-IID Settings
 To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios.
@@
 This phenomenon leads to Non-IID data across organizations.
-The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a certain percentage of all classes.
-However, in cross-silo scenarios, this contradicts the assumptions of Dirichlet distribution.
-To model the data distribution in cross-silo scenarios, we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts.
+The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a positive percentage of every class.
+However, in cross-silo scenarios, some clients may have no samples from certain classes, which contradicts the assumptions of Dirichlet distribution.
+To model the data distribution in cross-silo scenarios, similar to previous works [22][53], we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts.
 In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity.
@@
 ((c)) Avg=5, Std=2
-Fig. 2: Heat maps for different data distributions under heterogeneity distribution.
-7.1.4 Setting of Data Poisoning Attacks
+Figure 2: Heat maps for different data distributions under heterogeneity distribution.
+VII-A4 Setting of Data Poisoning Attacks
 In our experiments, we consider two types of data poisoning attacks: feature attacks and label attacks.
@@
 Formally, Att is set as the proportion of malicious clients.
-7.1.5 Evaluation Measure
+VII-A5 Evaluation Measure
 Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks.
@@
 For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks.
-In addition, we compare our PPFPL with robust schemes (i.e., Krum [44] and Foolsgold [45]) and privacy-preserving robust schemes (i.e., ShieldFL [18] and PBFL [16]) under the same experiment conditions.
-7.2 Experimental Results
-7.2.1 Visualization of Prototypes
-We use t-SNE to visualize client-submitted prototypes in PPFPL, as shown in Fig.
+In addition, we compare our PPFPL with robust schemes (i.e., Krum [35], Foolsgold [54], and FedDMC [36]) and privacy-preserving robust schemes (i.e., ShieldFL [18], PBFL [16], and AntidoteFL [20]) under the same experiment conditions.
+VII-B Experimental Results
+VII-B1 Visualization of Prototypes
+We visualize the prototypes submitted by clients in PPFPL using t-SNE, as shown in Fig.
 3.
-Specifically, we display prototypes from clients under feature and label attacks with Att =20% on the CIFAR10 dataset.
-This figure shows the distribution of these prototypes in a single communication round.
+Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on the CIFAR10 dataset within a single communication round.
 We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks.
@@
 This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas.
-To mitigate the severe impact of malicious clients, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results.
+Most importantly, we observe a clear difference between the prototypes submitted by benign and malicious clients.
+This difference is prominent because it is independent of the data distribution.
+The finding validates the effectiveness of using prototypes to defend against data poisoning attacks.
+To filter out these malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results.
 The following is the security evaluation of PPFPL.
-((a)) Feature attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
-((b)) Feature attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
-((c)) Label attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
-((d)) Label attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
-Fig.
-3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
-7.2.2 Security Evaluation
+((a)) Feature attack,$\textit{Avg}=3,\textit{Std}=2$.
+((b)) Feature attack,$\textit{Avg}=5,\textit{Std}=1$.
+((c)) Label attacks, $\textit{Avg}=3$, $\textit{Std}=2$.
+((d)) Label attacks, $\textit{Avg}=5$, $\textit{Std}=1$.
+Figure 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
+VII-B2 Security Evaluation
 ((a)) Feature attacks, FMNIST.
@@
 ((d)) Label attacks, CIFAR10.
-Fig. 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
+Figure 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
 ((a)) Feature attacks, CIFAR10.
 ((b)) Label attacks, CIFAR10.
-Fig. 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
+Figure 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
 To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$.
 Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$.
-Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, and PBFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
+Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, and AntidoteFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
 The experiment results are shown in Fig.
 4, where the black dashed line represents the highest accuracy of PPFPL without attacks.
-We observe that the accuracies of ShieldFL and PBFL (under attacks) are severely lower than FedAvg (no attack), especially for CIFAR10.
+We observe that the accuracy of ShieldFL and PBFL (under attacks) is severely lower than FedAvg (no attack) in CIFAR10.
 This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious.
 This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance.
-In contrast, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence.
+Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of the model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning.
+Moreover, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence.
 This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients.
 Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
-Additionally, we evaluate the change of loss during training for PPFPL and the existing schemes, as shown in Fig.
+Additionally, we evaluate the change of loss during training for PPFPL and the compared schemes, as shown in Fig.
 5.
@@
 Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security.
-Notably, our experiments show that the performance impact of feature attacks or label attacks is similar for PPFPL and other schemes.
+Notably, our experiments show that the performance impact of feature attacks and label attacks is similar for PPFPL and other schemes.
 This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
@@
 ((b)) Label attacks
-Fig. 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
-7.2.3 Different Data Distributions for PPFPL
+Figure 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
+VII-B3 Different Data Distributions for PPFPL
 To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10.
-In addition, we compare the performance of Krum, Foolsgold, ShieldFL, and PBFL.
-The specific experimental results are shown in Table 1.
-Notably, since test average accuracy has fluctuated after each communication round, we are select the average of five highest test average accuracies across the communication rounds.
-From Table 1, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution.
+In addition, we compare the performance of Krum, Foolsgold, ShieldFL, PBFL, FedDMC, and AntidoteFL.
+The specific experimental results are reported in Table I.
+Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across the communication rounds.
+From Table I, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution.
 This is attributed to that the client-submitted prototype does not change due to the change in data distribution.
-In addition, we observe a slight degradation in the performance of PPFPL with higher proportion of attacks.
+In addition, we observe a slight degradation in the performance of PPFPL with $Att=30\%$ attacks.
 This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon.
 Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
-Table 1: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
+TABLE I: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
 Dataset Method Att % Std Test Average Accuracy (%)
@@
 30 1 95.78 94.15 95.95 96.50 96.75
-2 95.47 93.62 93.48 94.70 94.62
-ShieldFL 20 1 96.91 94.05 95.14 95.64 95.25
+2 95.40 93.62 93.48 94.70 94.62
+ShieldFL 20 1 96.91 94.05 95.10 95.64 95.25
 2 97.25 94.88 95.69 95.44 95.76
 30 1 96.77 93.79 94.93 95.60 95.81
+2 97.08 94.69 95.46 95.73 94.46
+FedDMC 20 1 97.62 97.46 97.16 96.86 96.53
+2 97.53 97.32 97.04 96.73 96.39
+30 1 97.62 96.98 96.88 96.46 96.17
+2 97.43 96.58 96.34 96.26 94.33
+AntidoteFL 20 1 97.58 97.40 97.09 96.73 96.34
+2 97.31 97.18 96.85 95.45 96.30
+30 1 97.35 96.39 96.42 96.06 96.20
 2 97.08 94.69 95.46 95.73 94.46
@@
 2 75.81 77.35 70.57 66.42 65.46
+FedDMC 20 1 90.60 88.21 87.03 86.40 86.11
+2 89.53 88.70 87.27 87.20 87.13
+30 1 89.73 88.03 86.38 86.70 86.14
+2 89.40 87.52 86.23 86.40 86.12
+AntidoteFL 20 1 89.13 87.20 87.03 86.40 86.11
+2 88.72 86.13 85.34 85.16 86.48
+30 1 87.25 86.24 85.13 84.10 84.25
+2 87.10 86.61 85.10 84.39 84.90
 PPFPL 20 1 91.38 90.18 89.27 88.56 88.13
@@
 2 57.10 56.76 65.47 66.42 67.69
+FedDMC 20 1 75.28 74.16 74.37 73.86 73.50
+2 73.52 72.40 72.37 72.06 71.58
+30 1 74.60 74.28 73.97 73.62 73.39
+2 72.67 72.48 72.11 71.85 70.78
+AntidoteFL 20 1 72.66 72.27 71.19 72.19 70.65
+2 70.86 70.40 70.19 69.84 69.75
+30 1 71.49 71.61 70.43 71.24 69.20
+2 69.82 68.70 68.42 68.83 68.17
 PPFPL 20 1 83.41 80.95 77.82 74.76 71.35
@@
 2 83.16 82.37 77.47 74.62 71.19
-7.2.4 Secure Aggregation Protocol for PPFPL
+VII-B4 Secure Aggregation Protocol for PPFPL
 To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
-In addition, the $\boldsymbol{X}$ in Table 2 denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging.
-From Table 2, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings, but the magnitude of that degradation is limited.
-This is because each client minimizes its local classification loss and aligns local prototype with global prototype during local model training.
+In addition, the $\boldsymbol{X}$ in Table II denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging.
+As shown in Table II, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings.
+The performance degradation stems from the influence of malicious clients’ prototypes on the aggregation process, leading to a poisoned global prototype.
 However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss.
 Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients.
-Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which suggests that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol.
+Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol.
 The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks?
-Table 2: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
+We explore this question in the next experiment.
+TABLE II: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
 Datasets Att $\chi$ = -1 $\chi$ = 0 $\chi$ = 0.2 $\chi$ = 0.5 $\boldsymbol{X}$
@@
 ((d)) Att = 80%
-Fig. 7: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
-7.2.5 High Proportion Attacks for PPFPL
-To test the performance of PPFPL under high proportion attacks, we evaluate the performance of PPFPL under Att = 20%, 40%, 60% and 80%, respectively.
+Figure 7: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
+VII-B5 High Proportion Attacks for PPFPL
+To test the performance of PPFPL under high proportion attacks, we evaluate it with Att = 20%, 40%, 60% and 80%, respectively.
 In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
@@
 Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
-7.2.6 Stability of PPFPL
+VII-B6 Stability of PPFPL
 ((a)) Feature attacks
 ((b)) Label attacks
-Fig. 8: Stability of PPFPL under MNIST, FMNIST, and CIFAR10.
+Figure 8: Stability of PPFPL under MNIST, FMNIST, and CIFAR10.
 It is observed from Fig.
-4 and 5 that the performance of PPFPL fluctuates.
+4 and 5 that the performance of PPFPL exhibits some fluctuations.
 Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments.
-To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of participants being malicious.
+To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious.
 The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
@@
 The stability of PPFPL proves its reliability in real-world scenarios.
-7.2.7 Scalability of PPFPL
+VII-B7 Scalability of PPFPL
 Federated learning is often deployed with an uncertain number of clients, and scalability directly determines whether PPFPL can maintain stable performance in environments with different numbers of clients.
@@
 ((c)) CIFAR10
-Fig. 9: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
-7.2.8 Impact of Local Iterations
+Figure 9: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
+VII-B8 Impact of Local Iterations
 We investigate the effect of the number of local iterations on the test average accuracy for PPFPL.
@@
 ((c)) CIFAR10.
-Fig. 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
-7.2.9 Efficiency Evaluation
+Figure 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
+VII-B9 Efficiency Evaluation
 ((a))
 ((b))
-Fig.
-11: Communication parameters.
+Figure 11: Communication parameters.
 (a) The number of parameters submitted by clients under different local models.
 (b) The number of parameters submitted by client using CNN model in one round under different scheme.
-To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients in PPFPL, as shown in Fig.
+To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as shown in Fig.
 11.
-We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same parameters.
+We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same amount of parameters.
 Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture.
-This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, not the model architecture.
+This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, rather than the model architecture.
 Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
-Table 3: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
+TABLE III: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
        Operation        PBFL        ShieldFL        PPFPL
@@
        Decrypt        1.35s        1.19s        0.40s
-In addition, ciphertext operations are the most important factor affecting the efficiency of our framework.
+In addition, ciphertext operations are the primary factor affecting the efficiency of our framework.
 We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations.
@@
 For each client, we use the time spent on encryption and decryption for each iteration as a metric.
-As shown in Table 3, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL.
+As shown in Table III, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL.
 This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
-7.3 Complexity Analysis
-Table 4: Computation Overload and Communication Overload of client.
+VII-C Complexity Analysis
+TABLE IV: Computation Overload and Communication Overload of client.
 Scheme Computation Overload Communication Overload
@@
 We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15].
-The results are shown in Table 4.
+The results are shown in Table IV.
 The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption.
@@
 Compared with [16][18][15], the relationship $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds.
-8 Concludes
+VIII Concludes
 In this paper, we propose PPFPL, a privacy-preserving federated prototype learning framework, which improves performance while resisting data poisoning attacks in poisoned Non-IID data.
@@
 As for how to implement lightweight privacy-preserving FL to resist data poisoning attacks on Non-IID data, we leave them to future exploration.
-Acknowledgements
-This work was supported by the National Natural Science Foundation of China under Grants 62272256 and 62202250, the Major Program of Shandong Provincial Natural Science Foundation for the Fundamental Research under Grant ZR2022ZD03, the National Science Foundation of Shandong Province under Grant ZR2021QF079, the Talent Cultivation Promotion Program of Computer Science and Technology in Qilu University of Technology (Shandong Academy of Sciences) under Grant 2023PY059, the Pilot Project for Integrated Innovation of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences) under Grant 2022XD001, and the Colleges and Universities 20 Terms Foundation of Jinan City under Grant 202228093.
-CRediT authorship contribution statement
-Hongliang Zhang: Writing – original draft, Validation, Software, Methodology, Conceptualization.
-Jiguo Yu: Writing – review & editing, Supervision.
-Fenghua Xu: Software, Validation, Writing – review & editing.
-Chunqiang Hu: Writing – review & editing.
-Yongzhao Zhang: Writing – review & editing, Supervision.
-Xiaofen Wang: Software, Validation.
-Zhongyuan Yu: Validation.
-Xiaosong Zhang: Supervision.
-Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Data availability
-Data will be made available on request.

### v5

@@
-PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data
+PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks
 Abstract
-Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train by submitting hidden model updates.
-Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients in cross-silo scenarios.
-Existing solutions have struggled to improve the performance of PPFL in poisoned Non-IID data.
-To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while protecting the privacy of clients.
-Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distribution.
-In addition, we design a secure aggregation protocol by utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, which greatly reduces the impact of malicious clients.
-Theoretical analyses confirm the convergence and privacy of PPFPL, and experimental results on publicly datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID data.
-Index Terms:
+Privacy-Preserving Federated Learning (PPFL) enables multiple clients to collaboratively train models by submitting secreted model updates.
+Nonetheless, PPFL is vulnerable to data poisoning attacks due to its distributed training paradigm in cross-silo scenarios.
+Existing solutions have struggled to improve the performance of PPFL under poisoned Non-Independent and Identically Distributed (Non-IID) data.
+To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance against poisoned Non-IID data while protecting client privacy.
+Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distributions.
+In addition, we design a secure aggregation protocol utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, significantly reducing the impact of malicious clients.
+Theoretical analyses confirm the convergence and privacy of PPFPL.
+Experimental results on public datasets show that PPFPL effectively resists data poisoning attacks under Non-IID settings.
+{IEEEkeywords}
 Privacy-Preserving, Federated Learning, Cross-Silo, Data Poisoning Attacks, Poisoned Non-IID Data.
-I Introduction
+1 Introduction
 Federated Learning (FL) is a distributed learning paradigm where each client shares its model updates instead of raw training data.
-In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1].
+In industrial applications, large volumes of data are distributed across independent organizations governed by strict privacy regulations [1].
 To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3].
-Specifically, in cross-silo FL, the clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities.
-However, client-submitted model updates are vulnerable to privacy attacks, which threats the privacy security of cross-silo FL [4][5].
+Specifically, in cross-silo FL, clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities.
+However, client-submitted model updates are vulnerable to privacy attacks, threatening the security of cross-silo FL [4][5].
 To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] and Homomorphic Encryption (HE)-based [8][9] PPFL approaches are proposed.
-In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by introducing noise.
-Conversely, HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
-While HE-based approaches demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11].
-Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13].
-Moreover, Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and more difficult to defend.
-To audit obscured model updates, existing methods [14, 15, 16, 17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates in ciphertext.
+In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by injecting noise.
+Conversely, HE-based approach provides higher privacy security without sacrificing the global model accuracy, making it more suitable for cross-silo FL.
+While HE-based approaches have demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11].
+Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data and submitting model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13].
+Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and difficult to defend.
+To audit secreted model updates, existing methods [14, 15, 16, 17, 18] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates within ciphertext.
 However, these works overlook the data heterogeneity in FL, i.e., the data among clients is typically non-independent and identically distributed (Non-IID).
-The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17] to distinguish whether the deviations are caused by data heterogeneity or by malicious model updates.
-To distinguish between benign and malicious updates in the Non-IID data, existing methods [18, 19, 20] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase, thereby filtering out updates caused by model poisoning attacks.
-However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, thereby forming poisoned Non-IID data.
-These manipulations distort the optimization direction of local training, causing these methods [18, 19, 20] to optimize toward the tampered Non-IID distribution, thereby degrading the performance of FL.
-Therefore, a critical challenge remains: how to enhance the performance of PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
-Inspired by prototype learning [21], several works [22, 23, 24, 25] proposed prototype aggregation to address the Non-IID issue in FL, where prototypes are exchanged between the server and clients.
-Each prototype represents a class-level feature, computed as the mean of the feature representations of samples belonging to the same class.
-For instance, the work in [23] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”.
-By exchanging these prototypes, clients gain more knowledge about the concept of “cat”.
-Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
-[…]
+The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17, 18] to distinguish whether the deviations are caused by data heterogeneity or by malicious manipulation.
+To distinguish between benign and malicious updates under Non-IID data, existing methods [19, 20, 21] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase.
+However, in data poisoning attacks, malicious clients manipulate the features and labels of their training data, thereby forming poisoned Non-IID data.
+These manipulations distort the optimization direction of local training, causing these methods [19, 20, 21] to optimize toward the tampered Non-IID distribution, severely degrading the performance of FL.
+Therefore, a critical challenge remains: how to enhance the performance of PPFL under poisoned Non-IID data while effectively resisting data poisoning attacks.
+Inspired by prototype learning [22], several works [23, 24, 25, 26] have introduced prototype aggregation to address the Non-IID problem in FL by exchanging prototypes between servers and clients.
+Each prototype represents a class-level feature, computed as the mean of the feature representations of samples within the same class.
+For instance, the work in [25] suggests that in recognizing the class “cat,” different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”.
+By prototype exchange, clients gain more knowledge about the concept of “cat”.
+Consequently, prototypes, as feature representations independent of local data distributions, motivate our work.
+In cases of poisoned Non-IID data, both tampered data features and distributions cannot be repaired since the server cannot control malicious client behavior.
+This prompts us to pose a question: Is it possible to design a PPFL that leverages prototype learning, so that client-submitted model updates are affected only by tampered data features rather than by tampered data distribution, while incorporating a secure aggregation protocol to eliminate the impact of tampered data features, thereby achieving Byzantine-robust results (i.e., guaranteeing reliability under arbitrary and potentially malicious behaviors in distributed computing systems [27])?
+To answer this question, we propose a Privacy-Preserving Federated Prototype Learning framework, named PPFPL, which is suitable for cross-silo scenarios.
+The PPFPL framework consists of two “non-colluding” servers (assumed to be honest but curious) and multiple clients.
+The key novelty of PPFPL lies in: $(\textbf{\text{i}})$ We propose a novel local optimization function leveraging prototype learning, where each client submit prototypes to two servers, which mitigates the impact of tampered data distributions.
+Unlike existing prototype-based works [23][24], our optimization function employs cosine similarity to constrain the influence of malicious clients (see the proof of Theorem 1).
+$(\textbf{\text{ii}})$ We design a secure aggregation protocol between two servers to aggregate client-submitted prototypes using HE and SMC techniques.
+This protocol achieves Byzantine-robust results while preserving the privacy of benign clients.
 Our main contributions are summarized as follows.
-  • We are the first to introduce prototype learning into PPFL to defend against data poisoning attacks.
+  • To the best of our knowledge, this is the first work that introduces prototype learning into PPFL to defend against data poisoning attacks.
 By transferring prototypes between clients and servers, our framework mitigates the impact of tampered data distribution and enhances the FL performance in poisoned Non-IID data.
-  • We utilize HE and SMC techniques to design a secure aggregation protocol across two servers to filter prototypes submitted by malicious clients, ensuring Byzantine-robust results while protecting the privacy of benign clients.
-  • We provide a convergence guarantee for PPFPL in the presence of data poisoning attacks, which theoretically ensures the framework’s feasibility.
+  • We employ HE and SMC techniques to design a secure aggregation protocol across two servers that filters malicious prototypes, thereby ensuring Byzantine-robust results while protecting the privacy of benign clients.
+Notably, existing studies have not provided solutions for privacy protection in prototype-based federated learning, and our work fills this important research gap.
+  • We establish theoretical guarantees on convergence and privacy for PPFPL, thereby ensuring the framework’s feasibility.
   • Compared to existing methods, the superiority of our framework has been empirically validated in poisoned Non-IID data.
 The rest of the paper is organized as follows.
-Section II reviews PPFL works against poisoning attacks.
-Section III introduces prototype learning and HE.
-Section IV formalizes the system model.
-Our PPFPL is detailed in Section V.
-Section VI provide theoretical analysis.
-Section VII reviews experimental results.
-Finally, Section VIII concludes this paper.
-II RELATED WORK
-Privacy preservation in FL: Although FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients.
-To resist such attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates.
+Section 2 reviews PPFL works against poisoning attacks.
+Section 3 introduces prototype learning and HE.
+Section 4 formalizes both the system model and threat model.
+Our PPFPL is detailed in Section 5.
+Section 6 provides theoretical analysis.
+Section 7 reviews experimental results.
+Finally, Section 8 concludes this paper.
+2 RELATED WORK
+Privacy preservation in FL: Although FL inherently provides a degree of privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients.
+To resist such attacks, DP-based and HE-based approaches have been proposed to preserve client-submitted model updates.
 Specifically, DP-based schemes [28, 29, 30, 31] deploy local differential privacy into model updates, ensuring privacy without compromising the utility of model updates.
-Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which degrades the FL performance to some extent.
-In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy.
-Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [30].
+Despite their low computation overheads, these schemes introduce Gaussian noise or Laplace noise into local model training, which inevitably degrades the FL performance to some extent.
+In contrast, HE is a commonly adopted cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy.
+Fang and Qian are one of the first scholars to implement PPFL using HE [32].
 They proposed a multi-party machine learning scheme using Paillier [33] technique without compromising clients’ privacy.
 Considering the heavy communication overhead of Paillier, the work in [34] proposed a privacy-preserving FL using CKKS (i.e., Cheon-Kim-Kim-Song) that reduces computational overhead associated with ciphertexts.
-This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [33].
+This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models compared to Paillier [35].
 However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
 Resisting data poisoning attacks in FL: Tolpegin et al.
-demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [34].
-Additionally, they proposed an aggregated defense strategy that identifies malicious clients in FL to circumvent data poisoning attacks.
-Similarly, the works in [35, 36, 37] proposed detection methods to filter malicious model updates.
+demonstrated that data poisoning attacks can severely reduce the classification accuracy of FL, even with a small percentage of malicious clients [36].
+Additionally, they proposed a defense strategy during aggregation that identifies malicious clients to circumvent data poisoning attacks.
+Similarly, the works in [37, 38, 39] introduced detection mechanisms to filter malicious model updates.
 Differently, Doku et al.
-employed SVM to audit client’s local training data for excluding malicious clients [38].
+employed Support Vector Machine (SVM) to audit client’s local training data for excluding malicious clients [40].
 However, this work violates the privacy of clients to some extent.
 Furthermore, the threats of privacy and data poisoning attacks usually coexist in practice across-silo FL.
-PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while concealing data poisoning attacks from malicious clients.
-Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17] combine HE and SMC to identify anomalous model updates in ciphertexts.
-For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function.
-Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks.
-However, these schemes [15, 14, 16, 17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious model updates or Non-IID data.
+While PPFL approaches employ cryptographic primitives to ensure client privacy, they mask data poisoning attacks from malicious clients.
+Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17, 18] integrate HE and SMC to detect anomalous model updates directly within ciphertexts.
+For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers to punish malicious clients via gradient extraction of logarithmic function.
+Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers, which protects data privacy and adjust the weight of clients’ gradients to weaken data poisoning attacks.
+Notably, these works [15, 14, 16, 17, 18] adopt a two-server architecture instead of a single server, which facilitates secure aggregation in adversarial scenarios.
+However, they neglect the deviations caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious manipulation or Non-IID data.
 To audit malicious updates in PPFL with Non-IID data, the works in [19][20] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage.
-Specifically, ShieldFL designs a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [18].
+Specifically, ShieldFL designed a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [19].
 Furthermore, Chen et al.
-adopts clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [19].
+adopt clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [20].
 These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data.
 They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[41], FedDyn[42], or FedLC[43], etc.) to radically improve performance of FL.
-These specialized techniques adds an auxiliary term into local optimization function to constrain model updates of clients, which helps to increase consistency of model updates during local model training.
-However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs.
-Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
-Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue.
+These specialized techniques introduce auxiliary terms into local optimization function to constrain client model updates, thereby improving consistency of model updates during local model training.
+However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, generating poisoned Non-IID data.
+This manipulation misleads specialized training techniques to adjust model updates based on these compromised inputs, which ultimately cause the global model to converge to a malicious objective.
+Therefore, the above works [19][20] cannot effectively integrate these specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
+Federated prototype learning: Recently, prototype learning has been applied in federated learning to address the Non-IID issue.
 Specifically, the works in [23, 25, 24, 26] are one of the first to propose federated prototype learning using the concept of prototype learning.
-Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes.
+Different from the specialized training techniques (e.g., FedProx, FedDyn, and FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes.
 In other words, each class holds its corresponding prototype that is independent of other classes.
-As a result, client-submitted prototype is affected by the samples but is independent of data distribution among clients.
-This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
-Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks.
-Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation caused by tampered Non-IID data in across-silo FL, while preserving the privacy of client-submitted prototypes.
-III PRELIMINARIES
-This section introduces prototypes in federated learning and CKKS technology.
-III-A Prototypes Meet Federated Learning
+Consequently, client-submitted prototype is affected by the samples but is independent of data distribution among clients.
+This insight motivates our work.
+Although previous works [23, 25, 24, 26] have employed prototype learning to enhance the FL performance on Non-IID data, they are vulnerable to data poisoning attacks in distributed scenarios.
+Specifically, malicious clients can compromise the performance of federated training by uploading their poisoned prototypes.
+Furthermore, existing work [44] designed a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes, which seriously damages the client’s privacy.
+To this end, our proposed PPFPL mitigates the performance degradation caused by data poisoning attacks on Non-IID data while ensuring client privacy.
+3 PRELIMINARIES
+This section introduces prototypes in federated learning and CKKS technology. In addition, TABLE 1 provides explanations of the acronyms used in this paper.
+Table 1: Summary of main acronyms.
+Notation Meaning Notation Meaning
+FL Federated Learning PPFL Privacy-Preserving Federated Learning
+HE Homomorphic Encryption CKKS Cheon–Kim–Kim–Song encryption
+SMC Secure Multi-Party Computation DP Differential Privacy
+KGC Key Generation Center SVM Support Vector Machine
+Non-IID
+Non-Independent and
+Identically Distributed
+RSA
+Rivest–Shamir–Adleman cryptosystem
+3.1 Prototypes Meet Federated Learning
 In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [45].
-Due to this property, the prototypes of the same class among clients are similar in FL task.
-Consequently, many FL schemes [22, 23, 24, 25] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
-To further understand the prototype calculation in federated prototype learning, we introduce some basic notations below.
-Let $\mathcal{S}$ be the set of clients, where each client $m\in\mathcal{S}$ has an independent private dataset, denoted as $\mathcal{D}_{m}=\{(\textit{{x}}_{(i)},y_{(i)})\}^{|\mathcal{D}_{m}|}$.
-Here, $|\mathcal{D}_{m}|$ represents the number of samples in client $m$, and $(\textit{{x}}_{(i)},y_{(i)})$ denotes sample $i$ in dataset, where $\textit{{x}}_{(i)}$ and $y_{(i)}$ correspond to the feature vector and class label of sample $i$, respectively.
+This inherent property ensures that prototypes of the same class are similar in FL task.
+Consequently, many FL schemes [23, 25, 24, 26] employ prototype learning to address Non-IID challenges, enabling clients to align their local prototypes with global class representations during local model training.
+To understand the prototype calculation in federated prototype learning, we introduce some basic notations below.
+Let $\mathcal{S}$ be the set of clients, where each client $m\in\mathcal{S}$ owns a private dataset, denoted as $\mathcal{D}_{m}=\{(\textit{{x}}_{(i)},y_{(i)})\}^{|\mathcal{D}_{m}|}$ .
+Here, $|\mathcal{D}_{m}|$ is the number of samples in client $m$, and $(\textit{{x}}_{(i)},y_{(i)})$ denotes sample $i$ in dataset, where $\textit{{x}}_{(i)}$ and $y_{(i)}$ correspond to the feature vector and class label of sample $i$, respectively.
 Meanwhile, let $\mathcal{I}$ be the set of classes in classification task, where each class $k$ belongs to $\mathcal{I}$.
 In classification task, the local model includes a feature extractor and a decision classifier.
-Specifically, the feature extractor transforms sample features into compressed features, while decision classifier maps the compressed features to get classification results.
+Specifically, the feature extractor transforms raw sample features into compressed representations, while the decision classifier maps the compressed features to get classification results.
 Formally, let $f_{m}(\textit{{r}}_{m,t};\cdot)$ be feature extractor for client $m$, parameterized by $\textit{{r}}_{m,t}$, where $t$ denotes the $t$-th communication round.
-Given the feature $\textit{{x}}_{(i)}$ of sample $i$, it is input to feature extractor to obtain compressed feature $\textit{{u}}_{(i)}=f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)})$.
+Given the feature $\textit{{x}}_{(i)}$ of sample $i$, it is input into feature extractor to obtain compressed feature $\textit{{u}}_{(i)}=f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)})$.
 Let $g_{m}(\textit{{z}}_{m,t};\cdot)$ be decision classifier for client $m$, parameterized by $\textit{{z}}_{m,t}$.
 The classifier maps the compressed feature $\textit{{u}}_{(i)}$ to predict the class $y^{\prime}=g_{m}(\textit{{z}}_{m,t};\textit{{u}}_{(i)})$.
-Thus, we denote the local model as $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=g_{m}(\textit{{z}}_{m,t};\cdot)\circ f_{m}(\textit{{r}}_{m,t};\cdot)$, where $\circ$ denotes composite operator.
-To simplify notation, we use $\textit{{w}}_{m,t}$ to denote $(\textit{{r}}_{m,t},\textit{{z}}_{m,t})$, so we have $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=\mathcal{F}_{m}(\textit{{w}}_{m,t};\cdot)$, and $\textit{{w}}_{m,t}$ is consider as model parameters for client $m$.
-Next, we introduce calculation process of prototypes.
-In federated prototype learning, prototypes can be categorized into local prototypes, computed by clients, and global prototypes, aggregated by the server.
-Specifically, each client’s goal is to align its local prototype to global prototype during local model training.
-Thus, each client computes its local prototype via its training dataset during local model training.
-Formally, let $\textit{{c}}^{k}_{m,t}$ be the local prototype of class $k\in\mathcal{I}$ at client $m$ in $t$-th communication round, calculated as
-$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$ (1)
+Thus, we denote the local model as $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=g_{m}(\textit{{z}}_{m,t};\cdot)\circ f_{m}(\textit{{r}}_{m,t};\cdot)$, where $\circ$ denotes composition operator.
+For simplicity, we use $\textit{{w}}_{m,t}$ to denote $(\textit{{r}}_{m,t},\textit{{z}}_{m,t})$, so we have $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=\mathcal{F}_{m}(\textit{{w}}_{m,t};\cdot)$, and $\textit{{w}}_{m,t}$ is considered as model parameters for client $m$.
+Next, we present calculation process of prototypes.
+In federated prototype learning, prototypes can be categorized into local prototypes (computed by clients) and global prototypes (aggregated by the server).
+Specifically, each client’s goal is to align its local prototypes with global prototypes during local model training.
+Each client computes its local prototype via its training dataset during local model training.
+Formally, let $\textit{{c}}^{k}_{m,t}$ be the local prototype of class $k\in\mathcal{I}$ at client $m$ in $t$-th communication round, calculated as:
+$\begin{aligned} \textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},\end{aligned}$
 where $\mathcal{D}^{k}_{m}$ denotes the dataset with class $k$ at client $m$.
-The $\textit{{c}}^{k}_{m,t}$ can be understood as the mean of compressed features of samples belonging to class $k$ at client $m$.
-Further, we can get local prototypes for all classes at client $m$, denoted as $\{\textit{{c}}^{k}_{m,t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$.
-Once local model training is complete, each client submits its local prototypes to server.
-To calculate global prototype, the server adopts an averaging operation to local prototypes submitted by clients.
+$\textit{{c}}^{k}_{m,t}$ can be understood as the mean of compressed features of samples belonging to class $k$ at client $m$.
+Further, the set of local prototypes for all classes at client $m$ is denoted as $\{\textit{{c}}^{k}_{m,t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ .
+After completing local training, each client submits its local prototypes to the server for aggregation.
+To calculate global prototype, the server adopts an averaging operation on local prototypes submitted by clients.
 Thus, the global prototype $\textit{{C}}^{k}_{t+1}$ for each class is calculated as follows:
-$\textit{{C}}^{k}_{t+1}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\textit{{c}}_{m,t}^{k},\forall k\in\mathcal{I},$ (2)
+$\begin{aligned} \textit{{C}}^{k}_{t+1}&=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\textit{{c}}_{m,t}^{k},\forall k\in\mathcal{I},\end{aligned}$
 where $|\mathcal{S}|$ denotes the number of clients.
-For global prototypes of all classes, we denote them by the set $\{\textit{{C}}^{k}_{t+1}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$.
+The set of global prototypes for all classes is denoted as $\{\textit{{C}}^{k}_{t+1}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ .
 Subsequently, the server distributes the latest global prototypes to each client to further train their local model.
-III-B CKKS
-CKKS is a homomorphic encryption technology.
-Similar to traditional HE, it is a cryptographic technique based on mathematical computations, allowing operations on encrypted data without decrypting.
+3.2 CKKS Technique
+A ciphertext is the encrypted form of plaintext data, generated through an encryption algorithm.
+Homomorphic encryption allows operations to be performed directly on ciphertexts without decryption.
 However, traditional HE schemes such as RSA [46], ElGamal [47], and Paillier are limited in that they only support either additive or multiplicative operations but not both simultaneously.
 In contrast, the CKKS technology provides both additive and multiplicative homomorphic encryption, which is known as full HE [48].
-In addition, CKKS is known for its efficiency, especially in terms of encryption/decryption speed when dealing with scale vectors of varying parameter lengths.
-Therefore, we employ CKKS for the privacy protection of clients due to its computational efficiency.
-For a more detailed implementation principle and process of CKKS, please refer to paper [46].
-IV Problem Statement
+In addition, CKKS is known for its efficiency, especially in terms of encryption/decryption speed when applied to large-scale vectors with varying parameter lengths.
+Therefore, we employ CKKS to ensure client privacy while maintaining computational efficiency.
+The CKKS scheme mainly consists of key generation, encoding, encryption, addition, multiplication, decryption, and decoding.
+Here is a brief description of each operation:
+  1. 1.
+    Key generation.
+$\textit{KeyGen}(\cdot)$: Given a security parameter $\kappa$, the $\textit{KeyGen}(\cdot)$ generates secret key $Sk$, public key $Pk$ and evaluation key for ciphertext calculations.
+  2. 2.
+    Encoding. $\textit{Ecd}(\cdot)$: Given a $(\frac{N}{2})$-dimensional vector t and a scaling factor, the $\textit{Ecd}(\cdot)$ encodes the factor t into a polynomial $\mu$.
+  3. 3.
+    Encryption. $\textit{Enc}(\cdot)$: Given a polynomial $\mu$, the $\textit{Enc}(\cdot)$ encrypts $\mu$ by the public key $Pk$ and generates a ciphertext $\psi$.
+  4. 4.
+    Addition.
+$\textit{Add}(\cdot)$: Given a pair of ciphertexts $\psi_{1}$ and $\psi_{2}$, the $\textit{Add}(\cdot)$ outputs a ciphertext $\hat{\psi}=\psi_{1}\oplus\psi_{2}$, where $\hat{\psi}$ is the ciphertext of the sum of plaintexts of $\psi_{1}$ and $\psi_{2}$.
+  5. 5.
+[…]
+  6. 6.
+    Decryption. $\textit{Dec}(\cdot)$: Given a ciphertext $\psi$, the $\textit{Dec}(\cdot)$ generates a polynomial $\mu$ by the secret key $Sk$.
+  7. 7.
+    Decoding. $\textit{Dcd}(\cdot)$ : Given an input polynomial $\mu$ and a scaling factor $s$, the $\textit{Dcd}(\cdot)$ outputs the vector t.
+For more details on the implementation principles and procedures of CKKS, please refer to the work in [49].
+4 Problem Statement
 In the section, we formalize the PPFPL framework, define potential threats, and design goals.
-IV-A PPFPL framework
+4.1 PPFPL framework
 The framework of PPFPL consists of four entities, each of which has its specific function, as shown in Fig.
 1.
-The interaction among entities composes operation of whole system.
-The specific functions of each entity are outlined as follows.
+The interaction among these entities composes the operation of whole system.
+The roles of each entity are outlined as follows.
 Figure 1: The PPFPL framework.
@@
 ② After local training is completed, clients submit encrypted local prototypes to the Aggregator.
-③ Verifier and Aggregator perform secure aggregation protocol to get encrypted global prototypes.
-④ Aggregator distributes encrypted global prototypes to Clients.
+③ Verifier and Aggregator perform a secure aggregation protocol to get encrypted global prototypes.
+④ Aggregator distributes the encrypted global prototypes to Clients.
   • Key Generation Center (KGC).
-The entity is responsible for generating and managing keys of both Clients and Verifier, which are important elements to ensure security of encryption/decryption process.
+The entity is responsible for generating and managing the keys of both Clients and Verifier, which are essential elements to ensure security of encryption/decryption process.
   • Clients.
@@
 The aim of benign organizations is to get a better model by federated training.
-They has a pair of public/secret keys generated by KGC, denoted as $Pkx$/$Skx$.
+They have a pair of public/secret keys generated by KGC, denoted as PkxPkx/SkxSkx.
   • Aggregator. The Aggregator is a central server responsible for aggregating local prototypes submitted by clients.
   • Verifier.
-The Verifier is a non-colluding central server and calculates cooperatively with Aggregator to aggregate local prototypes.
-It has a pair of public/private keys generated by KGC, denoted as $Pkv$/$Skv$.
+The Verifier is a non-colluding central server that cooperates with the Aggregator to aggregate local prototypes.
+It has a pair of public/private keys generated by KGC, denoted as PkvPkv/SkvSkv.
 In our framework, we not only define the function of each entity, but also define potential threats.
-IV-B Potential Threats for PPFPL
+4.2 Potential Threats for PPFPL
 We discuss potential threats of PPFPL in detail.
 i) The KGC is a trusted institution (e.g., government, union).
-ii) The Aggregator and Verifier are considered as “non-colluding” and “curious but honest”.
+ii) The Aggregator and Verifier are assumed to be non-colluding and curious but honest.
 Specifically, we assume that Aggregator and Verifier do not collude to attack PPFPL.
-The assumption is reasonable in practice, since it is usually impossible for two well-known service providers to collude with each other due to legal regulations, company reputation [47].
+The assumption is reasonable in practice, since it is usually impossible for two well-known service providers to collude with each other due to legal regulations and company reputation [50].
 Furthermore, we assume that they follow the system’s protocol but may attempt to get sensitive information (i.e., raw training data) by inferring client-submitted local prototypes.
-Notably, although prototype is different from gradient, traditional inference attack methods cannot recover client’s private training data from prototype.
-However, the work in [48] designs a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes.
+Notably, although prototypes differ from gradients, traditional inference attack methods cannot recover client’s private training data from prototype.
+Although prototypes differ from gradients, and traditional inference attacks cannot directly reconstruct private training data from them, the work in [44] designs a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes.
 Thus, preserving the privacy of prototypes is necessary.
-In addition, we do not consider security attacks resulting from employee insider threats (e.g., compromised employees within the Aggregator or Verifier).
+Notably, we do not consider security attacks resulting from employee insider threats (e.g., compromised employees within the Aggregator or Verifier).
 iii) In practice, we cannot guarantee that all clients in the system are honest.
@@
 Specifically, benign clients are “honest but curious”.
-For malicious clients, they can collude together to infer sensitive information about benign clients.
-We consider that the proportion of malicious clients is less than 50%, which is a more realistic threat in cross-silo FL due to the high reputation of large organizations participating.
-Furthermore, malicious clients can launch data poisoning attacks, and submit malicious local prototypes derived from the poisoned data.
+For malicious clients, they can collude to infer sensitive information about benign clients.
+We assume that the proportion of malicious clients is less than 50%, which is a more realistic threat in cross-silo FL due to the high reputation of large organizations participating.
+Furthermore, malicious clients can launch data poisoning attacks by submitting malicious prototypes derived from poisoned data.
 In federated prototype learning, we define the following two representative types of data poisoning attacks from the perspective of data features and labels.
@@
 The malicious clients tamper with the labels of their training samples into other random labels to form the tampered data distribution, aiming to reduce the performance of the entire federated learning.
-IV-C Design Goals of PPFPL
-The goal of our study is to improve performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks.
+4.3 Design Goals of PPFPL
+The goal of our study is to improve model performance in cross-silo federated learning under poisoned Non-IID data while resisting data poisoning attacks.
 Specifically, we design the PPFPL framework to fulfill the following goals:
   • Security.
-PPFPL should guarantee the correct model training in the presence of data poisoning attacks with Non-ID data.
+PPFPL should guarantee the robust model training in the presence of data poisoning attacks with Non-IID data.
 In other words, the model performance of each benign client is not affected by data poisoning attacks under different data distributions.
-  • Privacy. PPFPL should ensure privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
-  • Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar frameworks, thereby reducing privacy computation overheads and communication overheads.
-V Design of PPFPL
+  • Privacy. PPFPL should ensure the privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
+  • Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar schemes, thereby reducing both privacy computation and communication overheads.
+5 Design of PPFPL
 In this section, we first provide the overview of PPFPL, and then describe each step in detail.
-V-A Overview of PPFPL
+5.1 Overview of PPFPL
 The execution process of PPFPL is summarized in Algorithm 1.
@@
 Specifically, PPFPL iteratively performs the following two steps:
-  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
+  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes before submitting them to Aggregator.
   • Step II.
-Secure Aggregation Protocol: The two servers verify the normalization of encrypted local prototypes submitted by clients, and performs the secure two-party computation to get encrypted global prototypes, and distribute them to each client.
-The above process is stopped until configured number of communication rounds $T$. In the following, we describe the process of Steps I and II in detail.
+Secure Aggregation Protocol: The two servers verify the normalization of encrypted local prototypes submitted by clients, and perform the secure two-party computation to get encrypted global prototypes, which are then distributed to each client.
+The above process repeats until configured number of communication rounds $T$. In the following, we describe the process of Steps I and II in detail.
 Input: $\mathcal{S}$, $\textit{{w}}^{\textit{init}}$, $E$, $T$.
@@
 Algorithm 1 Overview of PPFPL
-V-B Local Computation
-The local computation step includes two essential stages: local model training and prototype handling stages. The detail is outlined in Algorithm 2.
+5.2 Local Computation
+The local computation step includes two essential stages: local model training and prototype handling stages. The details are outlined in Algorithm 2.
 Input: $\mathcal{S}$, $\mathcal{D}_{m}$, $\eta$ , $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$, $E$.
 Output: Encrypted local prototypes.
-1 for each client $m$in $\mathcal{S}$ do
+1 for each client $m$ $\in$ $\mathcal{S}$ do
 2  Get global prototypes $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ from Aggregator;
-3       Decrypt $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ using its $Skx$;
+3  Decrypt $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ using its SkxSkx;
 4  $\textit{{w}}_{m,t}^{(E)}\leftarrow\textbf{Training}$($\textit{{w}}_{m,t-1}^{(E)}$, $\mathcal{D}_{m}$, $\eta$, $E$, $\{\textit{{C}}^{k}_{t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$);
 5  $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}\leftarrow\textbf{Handling}(\textit{{w}}_{m,t}^{(E)},\mathcal{D}_{m})$;
-6       Send $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator;
+6  Send
+$\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
+to Aggregator;
 7
@@
 Algorithm 2 Local Computation
-V-B1 Local Model Training
-During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype.
+5.2.1 Local Model Training
+During local model training, each client aims to minimize its classification loss while aligning its local prototype close to the global prototype.
 To achieve this, we design an auxiliary term by leveraging prototype learning in the local optimization function.
 Formally, the local optimization function of client $m$ is defined as:
-$\mathcal{L}(\textit{{w}}_{m};\mathcal{D}$ ,${}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})$ (3)
-$+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k}),\forall(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}_{m},\forall k\in\mathcal{I},$
-where $\mathcal{L_{S}}(\cdot,\cdot)$ is the classification loss function (e.g., cross-entropy loss function), $\lambda$ is the importance weight of auxiliary term, and $\mathcal{L_{R}}(\cdot,\cdot)$ is the auxiliary term, defined as
-$\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\frac{1}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}(1-\textit{sim}(\textit{{c}}_{m}^{k},\textit{{C}}^{k})),$ (4)
+$\begin{aligned} \mathcal{L}(\textit{{w}}_{m};\mathcal{D}&{}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})\\ &+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k}),\forall(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}_{m},\forall k\in\mathcal{I},\end{aligned}$
+where $\mathcal{L_{S}}(\cdot,\cdot)$ is the classification loss function (e.g., cross-entropy loss function), $\lambda$ is the importance weight of the auxiliary term, and $\mathcal{L_{R}}(\cdot,\cdot)$ is the auxiliary term, defined as:
+$\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\frac{1}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}(1-\textit{sim}(\textit{{c}}_{m}^{k},\textit{{C}}^{k})),$
 where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors.
-The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions.
-The optimization function ensures that each client reduces the classification loss while aligning its local prototype with the global prototype in direction.
+The cosine similarity ranges from 1-1 to $1$, where the value closer to 1``1" indicates similar vector directions, while value closer to 1``-1" means opposing directions.
+The optimization function ensures that each client reduces its classification loss while aligning its local prototype with the global prototype in direction.
 Next, we present the process of local model training.
-Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model.
-This iterative process is executed in Algorithm 3.
+Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model, as detailed in Algorithm 3.
 The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$.
-Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round.
+Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the (t1)(t-1)-th round as the starting point in the current $t$-th round.
 Subsequently, each client iteratively performs the following stages.
-  • In $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
+  • In the $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
   • The client inputs the training data $\mathcal{D}_{m}^{(e)}$ into the local model’s feature extractor to compute the local prototype $\textit{{c}}^{k,(e)}_{m,t}$ via formula (1).
@@
     $\textit{{g}}_{m,t}^{(e-1)}=\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e-1)};\mathcal{D}_{m}^{(e)},\textit{{c}}^{k,(e)}_{m,t},\textit{{C}}^{k}_{t}),$
-    where $\nabla$ denotes derivation operation. Then, the local model $\textit{{w}}_{m,t}^{(e-1)}$ is updated via the following formula:
+    where $\nabla$ denotes derivation operation. The local model parameters $\textit{{w}}_{m,t}^{(e-1)}$ are then updated as:
     $\textit{{w}}_{m,t}^{(e)}=\textit{{w}}_{m,t}^{(e-1)}-\eta\textit{{g}}_{m,t}^{(e-1)},$
     where $\eta$ is the local learning rate.
-After $E$ local iterations of the above process, each client derives the updated local model parameters $\textit{{w}}_{m,t}^{(E)}$.
+After completing $E$ local iterations, each client derives the updated local model parameters $\textit{{w}}_{m,t}^{(E)}$.
 Input: $\textit{{w}}_{m,t-1}^{(E)}$, $\mathcal{D}_{m}$, $\eta$, $E$, $\{\textit{{C}}^{k}_{t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$.
@@
 1 $\textit{{w}}_{m,t}^{(0)}=\textit{{w}}_{m,t-1}^{(E)}$;
-2 for each local iteration $e$in $\{1,2,\cdots,E\}$ do
+2 for each local iteration $e$ $\in$ $\{1,2,\cdots,E\}$ do
 3  Randomly sample $\mathcal{D}_{m}^{(e)}\subset\mathcal{D}_{m}$;
-4       for each class $k$in $\mathcal{I}$ do
+4  for each class $k$ $\in$ $\mathcal{I}$ do
 5     Calculate $\textit{{c}}^{k,(e)}_{m,t}$ from $\mathcal{D}^{k,(e)}_{m}$ with formula (1);
@@
 Algorithm 3 Training
-V-B2 Prototype Handling
+5.2.2 Prototype Handling
 The prototype handling stage consists of three key phases: prototype generation, normalization, and encryption, as illustrated in Algorithm 4.
@@
 To submit more representative prototypes, each client regenerates them via the local model parameters $\textit{{w}}_{m,t}^{(E)}$ .
-Formally, let $\textit{{c}}^{k}_{m,t}$ be the submitted local prototype, is computed as:
-$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$
+Formally, let $\textit{{c}}^{k}_{m,t}$ be the submitted local prototype, which is computed as:
+$\begin{aligned} \textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},\end{aligned}$
 where $\textit{{r}}_{m,t}^{(E)}$ is the parameters of feature extractor.
 Normalization
-Considering that malicious clients amplify their submitted local prototypes, the local prototype of each class in each client is normalized.
-Formally, the local prototype is normalized by the following formula:
+Considering that malicious clients amplify their submitted local prototypes, the local prototype of each class at each client is normalized. Formally, the normalized local prototype is given by:
 $\widetilde{\textit{{c}}}^{k}_{m,t}=\textit{{c}}^{k}_{m,t}/\|\textit{{c}}^{k}_{m,t}\|,\forall m\in\mathcal{S},k\in\mathcal{I},$
-where $\widetilde{\textit{{c}}}^{k}_{m,t}$ is a unit vector. After normalization, each local prototype needs to be encrypted for privacy protection.
+where $\widetilde{\textit{{c}}}^{k}_{m,t}$ is a unit vector. After normalization, each local prototype is encrypted to ensure privacy protection.
 Input: $\textit{{r}}_{m,t}^{(E)}$, $\mathcal{D}_{m}$
 Output: $\{\llbracket\textit{{c}}_{m,t}^{k}\rrbracket\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
-1 for each class $k$in $\mathcal{I}$ do
+1 for each class $k$ $\in$ $\mathcal{I}$ do
 2  $\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}^{|\mathcal{D}^{k}_{m}|}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)})$;
 3  $\widetilde{\textit{{c}}}^{k}_{m}=\textit{{c}}^{k}_{m,t}/\|\textit{{c}}^{k}_{m,t}\|$;
-4       Encrypt $\widetilde{\textit{{c}}}^{k}_{m,t}$ by Verifier’s $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$;
+4  Encrypt $\widetilde{\textit{{c}}}^{k}_{m,t}$ by Verifier’s PkvPkv to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$;
 5
@@
 Encryption
-To protect privacy of clients, they encrypts their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
-V-C Secure Aggregation
+To protect privacy of clients, they encrypt their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ , and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
+5.3 Secure Aggregation Protocol
 To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to filter malicious local prototypes and obtain Byzantine-robust global prototypes.
 The protocol consists of normalization verification and secure two-party computation.
-V-C1 Normalization verification
-Since malicious clients may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized.
+5.3.1 Normalization Verification
+Since malicious clients may amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized.
 Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product.
 Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1.
-If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$.
+If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype is not normalized, client $m$ is removed from the set of clients $\mathcal{S}$.
 After validation, Verifier sends the remaining client set $\mathcal{S}$ to Aggregator.
-V-C2 Secure Two-party Computation
-To achieve privacy preservation and Byzantine-robust aggregation results, two servers perform secure two-party computation to calculate global prototypes.
-The computing of global prototype is shown in Algorithm 5 and described as follows.
+5.3.2 Secure Two-party Computation
+To ensure privacy preservation and Byzantine-robust aggregation, two servers perform secure two-party computation to calculate global prototypes.
+The computation of the global prototypes is shown in Algorithm 5 and described as follows.
 Input: $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, $\mathcal{S}$, $\chi$.
@@
 9Verifier:
-10      Decrypt $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ with $Skv$;
+10 Decrypt $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ with SkvSkv;
 11  Compute $j^{k}_{m,t}$ by formula (8);
 12  Compute $\textit{Sum}^{k}_{t}$;
-[…]
+13  Encrypt $j^{k}_{m,t}$, $\textit{{V}}^{n}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using ClientsPkxPkx;
 14  Send $\textit{Sum}^{k}_{t}$, $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$, and $\llbracket\textit{{V}}^{n}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to Aggregator;
@@
 To this end, Aggregator computes the trusted prototype via the formula:
-$\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\llbracket\textit{{c}}_{m,t}^{k}\rrbracket_{Pkv},\forall k\in\mathcal{I},$ (5)
+$\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\llbracket\textit{{c}}_{m,t}^{k}\rrbracket_{Pkv},\forall k\in\mathcal{I},$
 where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round.
-$\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: 1) In terms of magnitude, this is because the normalized malicious prototype do not affect the magnitude of trusted prototype, only the direction of trusted prototype; 2) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients.
+$\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: (i) In terms of magnitude, this is because the normalized malicious prototypes do not affect the magnitude of trusted prototype, only the direction of trusted prototype.
+(ii) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients.
 Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation.
 To achieve the plaintext, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result.
-Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype.
-If the direction of local prototype is similar to that of trusted prototype, its credibility is higher.
-Formally, Aggregator computes the credibility of local prototype by the following formula,
-$\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ $=\llbracket\textit{sim}(\widetilde{\textit{{c}}}^{k}_{m,t},\textit{{C}}^{\prime k}_{t+1})\rrbracket_{Pkv}=\left\llbracket\frac{\widetilde{\textit{{c}}}^{k}_{m,t}\cdot\textit{{C}}^{\prime k}_{t+1}}{\|\widetilde{\textit{{c}}}^{k}_{m,t}\|\|\textit{{C}}^{\prime k}_{t+1}\|}\right\rrbracket_{Pkv}$ (6)
-$=\frac{1}{\|\textit{{C}}^{\prime k}_{t+1}\|}\left(\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\right),$
+Then, cosine similarity, a widely used metric for measuring the angle between two vectors, is employed to evaluate the credibility of local prototype.
+If the direction of local prototype is similar to that of trusted prototype, it is assigned higher credibility.
+Formally, Aggregator computes the credibility of local prototype by the following formula:
+$\begin{aligned} \llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}&=\llbracket\textit{sim}(\widetilde{\textit{{c}}}^{k}_{m,t},\textit{{C}}^{\prime k}_{t+1})\rrbracket_{Pkv}=\left\llbracket\frac{\widetilde{\textit{{c}}}^{k}_{m,t}\cdot\textit{{C}}^{\prime k}_{t+1}}{\|\widetilde{\textit{{c}}}^{k}_{m,t}\|\|\textit{{C}}^{\prime k}_{t+1}\|}\right\rrbracket_{Pkv}\\ &=\frac{1}{\|\textit{{C}}^{\prime k}_{t+1}\|}\left(\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\right),\end{aligned}$
 where $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ denotes the credibility of local prototype.
-Additionally, we set the detection threshold $\chi$ and define local prototypes with its credibility less than $\chi$ as anomalous local prototypes.
+Additionally, we define a detection threshold $\chi$ and classify local prototypes with credibility less than $\chi$ as anomalous local prototypes.
 During global prototype aggregation, we set the aggregation weight of anomalous local prototype to 0.
@@
 where $j^{k}_{m,t}$ denotes the aggregation weight of local prototype for class $k$ at client $m$ in the $t$-th communication round.
-However, Aggregator cannot directly compare $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ with $\chi$ under encryption domain.
+However, Aggregator cannot directly compare $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ with $\chi$ within the encryption domain.
 To implement comparison within encryption domain, the work in [51] offers a ciphertext comparison method that outputs the maximum value of homomorphic ciphertexts corresponding to two plaintexts in the range (0, 1) without decryption.
-However, since the cosine similarity ranges from -1 to 1, and the method cannot be applied directly.
+However, since cosine similarity ranges from -1 to 1, the method cannot be applied directly.
 To resolve the range mismatch, Aggregator sets $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket=\frac{1}{2}(\llbracket\textit{sim}^{k}_{m,t}\rrbracket+\llbracket 1\rrbracket)$ and $\chi^{\prime}=\frac{1}{2}(\chi+1)$ as input to the comparison method [51].
 This is because the size relationship between $\llbracket\textit{sim}^{k}_{m,t}\rrbracket$ and $\chi$ remains unchanged.
-Then, Aggregator encrypts $\chi^{\prime}$ to get $\llbracket\chi^{\prime}\rrbracket$, and calculates the maximum value $\llbracket h^{k}_{m,t}\rrbracket$ between $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $\llbracket\chi^{\prime}\rrbracket$ by Algorithm 6.
+Then, Aggregator employs Algorithm 6 to calculate the maximum value $\llbracket h^{k}_{m,t}\rrbracket$ between $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $\llbracket\chi^{\prime}\rrbracket$.
+Notably, in the Algorithm 6, $a$ and $b$ are defined as $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $b=\llbracket\chi^{\prime}\rrbracket$.
+From these definitions, $q_{1}=\frac{a+b}{2}$ and $q_{2}=\frac{a-b}{2}$ are directly computed.
+The parameter $d$ in Algorithm 6 is a hyper-parameter.
+The larger the value of $d$, the more accurate the output.
+Based on previous work [51], setting $d=29$ is sufficient to ensure that the output matches the top-16 bits coincide with those of the true maximum value.
+In addition, the design principle process of Algorithm 6 is detailed in works [51][52].
 Input: $(\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,\llbracket\chi^{\prime}\rrbracket)\in(\llbracket 0\rrbracket,\llbracket 1\rrbracket),d\in\mathbb{N}$
 Output: an max value of $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ or $\llbracket\chi^{\prime}\rrbracket$
-1 $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,b=\llbracket\chi^{\prime}\rrbracket$;
+1 Initialize $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,b=\llbracket\chi^{\prime}\rrbracket$;
 2 $q_{1}=\frac{(a+b)}{2},q_{2}=\frac{(a-b)}{2}$;
 3 $a_{0}=q_{2}^{2},b_{0}=q_{2}^{2}-1$;
-4 for each $n$in $(0,d-1)$ do
+4 for each $n$ $\in$ (0,d1)(0,d-1) do
 5  $a_{n+1}=a_{n}(1-\frac{b_{n}}{2})$;
@@
 Moreover, $\textit{min}^{k}_{t}$ denotes the smallest value for class $k$ in set $\{\textit{Round}(p\times h^{k}_{m,t},6)\}_{m\in\mathcal{S}}^{|\mathcal{S}|}$ .
-The $\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}$ means that the local prototype $\textit{{c}}^{k}_{m,t}$ satisfies $\textit{sim}^{k}_{m,t}<\chi$, thus its aggregation weight $j^{k}_{m,t}$ is 0.
+The condition $\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}$ means that the local prototype $\textit{{c}}^{k}_{m,t}$ satisfies $\textit{sim}^{k}_{m,t}<\chi$, thus its aggregation weight $j^{k}_{m,t}$ is 0.
 Additionally, Verifier calculates $\textit{Sum}^{k}_{t}=\sum_{m\in\mathcal{S}}j^{k}_{m,t}$ and encrypts $j^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using the client’s public key $Pkx$ to get $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$ and $\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ , and sends them to Aggregator.
-The Aggregator computes $\frac{1}{\textit{{V}}}\odot\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$, and aggregates the encrypted local prototypes to get the encrypted global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by the following formula,
+The Aggregator computes $\frac{1}{\textit{{V}}}\odot\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ , and aggregates the encrypted local prototypes to get the encrypted global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by the following formula:
 $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ $=\frac{1}{\textit{Sum}^{k}_{t}}\sum_{m\in\mathcal{S}}\llbracket j^{k}_{m,t}\rrbracket_{Pkx}\times\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}.$ (9)
@@
 After the above process is completed, FL executes the next communication round until a predefined number of rounds is reached.
-VI Analysis
+6 Analysis
 In the section, we provide both convergence analysis and privacy analysis for PPFPL.
-Specifically, we make the following assumptions similar to existing general frameworks [22][51] for loss function (3).
+Specifically, we make the following assumptions similar to existing general frameworks [23][54] for the local optimization function (3).
 Assumption 1.
@@
 $\begin{gathered}\left\|\nabla\mathcal{L}_{m,t}^{(e_{1})}-\nabla\mathcal{L}_{m,t}^{(e_{2})}\right\|_{2}\leq L_{1}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2},\end{gathered}$
-where $\mathcal{L}_{m,t}^{(e_{1})}$ denotes loss function at the $(tE+e_{1})$-th local iteration in client $m$. This implies the following quadratic bound,
-$\mathcal{L}_{m,t}^{(e_{1})}$ $-\mathcal{L}_{m,t}^{(e_{2})}$
-$\leq\left\langle\nabla\mathcal{L}_{m,t}^{(e_{2})},\left(\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2}^{2}.$
+where $\mathcal{L}_{m,t}^{(e_{1})}$ denotes loss function at the $(tE+e_{1})$-th local iteration in client $m$. This implies the quadratic bound:
+$\begin{aligned} \mathcal{L}_{m,t}^{(e_{1})}&-\mathcal{L}_{m,t}^{(e_{2})}\leq\left\langle\nabla\mathcal{L}_{m,t}^{(e_{2})},\left(\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2}^{2}.\end{aligned}$
 Assumption 2.
@@
 $\mathbb{E}_{\mathcal{D}_{m}^{(e)}\sim\mathcal{D}_{m}}\left[\textit{{g}}_{m,t}^{(e)}\right]=\nabla\mathcal{L}\left(\textit{{w}}_{m,t}^{(e)};\mathcal{D}_{m}^{(e)}\right)=\nabla\mathcal{L}_{m,t}^{(e)},$
-and its variance is bounded by $\sigma^{2}$:
-$\mathbb{E}[\|\textit{{g}}_{m,t}^{(e)}-\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e)})\|_{2}^{2}]\leq\sigma^{2}.$
-Based on the above assumptions, we have the following theorem and corollaries.
+and its variance is bounded by $\sigma^{2}$: $\mathbb{E}[\|\textit{{g}}_{m,t}^{(e)}-\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e)})\|_{2}^{2}]\leq\sigma^{2}$ .
+Based on the above assumptions, we formulate the following theorem and corollaries.
 Notably, we add “ $\frac{1}{2}$ ” into the local iteration, denoted as $\{\frac{1}{2},1,\cdots,E\}$ in our analysis.
-For example, $tE$ denotes the time step before local prototype aggregation, and $tE+\frac{1}{2}$ denotes the time step between local prototype aggregation and the first local iteration of the $t$-th round.
+For example, $tE$ denotes the time step before local prototype aggregation, and $tE+\frac{1}{2}$ denotes the time step between local prototype aggregation and the first local iteration in the $t$-th round.
 Theorem 1.
-In PPFPL, regardless of the percentage of malicious clients, for the $t$-th communication round, the variation of loss function for each benign client can be bounded as,
-$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
-where
-$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
+In PPFPL, under any proportion of malicious clients, for the $t$-th communication round, the variation of loss function for each benign client can be bounded as:
+$\begin{aligned} \mathbb{E}&\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),\end{aligned}$
+where $\begin{aligned} G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.\end{aligned}$
+We can observe that each variable in this upper bound is independent of the malicious clients, indicating that the impact of malicious clients is constrained.
+This constraint arises from incorporating the cosine similarity in the auxiliary term of our local optimization function, which restricts the update direction of the local model.
 Proof.
 Assuming that Assumption 1 holds, we can get
-$\mathcal{L}_{m,t}^{(1)}$ $\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}+\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\left(\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}$ (10)
 […]
 where (a) follows from $\textit{{w}}_{m,t}^{(1)}=\textit{{w}}_{m,t}^{(\frac{1}{2})}-\eta\textit{{g}}_{m,t}^{(\frac{1}{2})}$ . Taking expectation on both sides of formula (LABEL:102218311), we can get
-$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-\eta\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\rangle\right]+\frac{\eta^{2}L_{1}}{2}\mathbb{E}\left[\left\|\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}\right]$
 […]
-[…]
-[…]
-$=\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\|\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})}\|_{2}^{2}+\frac{\eta^{2}L_{1}}{2}Var(\textit{{g}}_{m,t}^{(\frac{1}{2})})$
-[…]
-where (b), (d) and (e) follow from Assumption 2, (c) follows from $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$.
-Then, during the local computation step, the loss function is updated $E$ times, and it can be bounded as:
-$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}.\$ (11)
+where (b), (d) and (e) are derived from Assumption 2, (c) follows from $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$ .
+During the local computation step, the loss function is updated $E$ times, and it can be bounded as:
+$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}.\\$
 Additionally, since a single communication round involves both local computation and secure aggregation, we need to compute the impact of the aggregation result for loss function of each benign client.
 Specifically, the loss function of each benign client at the $((t+1)E+\frac{1}{2})$ time step is represented as follows:
-$\mathcal{L}_{m,t+1}^{\frac{1}{2}}=\mathcal{L}_{m,t}^{E}+\mathcal{L}_{m,t+1}^{\frac{1}{2}}-\mathcal{L}_{m,t}^{E}$ (12)
-$\stackrel{{\scriptstyle}}{{=}}\mathcal{L}_{m,t}^{E}+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+2})-\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+1})$
-$\overset{}{=}\mathcal{L}_{m,t}^{E}-\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+2})+\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+1})$
-[…]
-where (f) follows from $-1\leq\textit{sim}(\cdot,\cdot)\leq 1$. Taking expectation on both sides of formula (12), we can get
-$\mathbb{E}\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]$ $\leq\mathcal{L}_{m,t}^{E}+2\lambda.$ (13)
-Thus, during the $t$-th communication round, according to the formula (11) and formula (13), the variation of loss function for each benign client can be bounded as,
-$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
-where
-$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
-Thus, Theorem 1 is proved. ∎
+$\begin{aligned} &\mathcal{L}_{m,t+1}^{\frac{1}{2}}=\mathcal{L}_{m,t}^{E}+\mathcal{L}_{m,t+1}^{\frac{1}{2}}-\mathcal{L}_{m,t}^{E}\\ &\stackrel{{\scriptstyle}}{{=}}\mathcal{L}_{m,t}^{E}+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+2})-\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+1})\\ &\overset{}{=}\mathcal{L}_{m,t}^{E}-\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+2})+\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+1})\\ &\overset{\text{(f)}}{\leq}\mathcal{L}_{m,t}^{E}+2\lambda,\end{aligned}$
+where (f) follows from $1\leq\textit{sim}(\cdot,\cdot)\leq 1$ .
+Notably, although different proportions of malicious clients can influence the direction of $\textit{{C}}^{k}_{t+1}$, the contribution of malicious prototypes is inherently bounded by the cosine similarity $\textit{sim}(\cdot,\cdot)$.
+In other words, regardless of the proportion of the malicious clients, PPFPL suppresses the impact of malicious clients, thereby maintaining the robustness of the aggregated results.
+Taking expectation on both sides of formula (12), we can get:
+$\begin{aligned} \mathbb{E}\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]&\leq\mathcal{L}_{m,t}^{E}+2\lambda.\end{aligned}$
+Thus, during the $t$-th communication round, according to the formula (11) and formula (13), the variation of loss function for each benign client can be bounded as:
+$\begin{aligned} \mathbb{E}&\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),\end{aligned}$
+where $G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda$ .
+Thus, Theorem 1 is proved.
+∎
 Corollary 1.
-Given any fixed $\lambda$ and $E$, the $G(\eta)$ is a convex function with respect to $\eta$.
+Given any fixed $\lambda$ and $E$, the $G(\eta)$ is convex with respect to $\eta$.
 Proof.
-To prove that $G(\eta)$ is a convex function for $\eta$, we need to prove that the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative. Thus, we have
-$\frac{dG(\eta)}{d\eta}=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2},$
+To prove that $G(\eta)$ is convex for $\eta$, we need to prove that the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative. Thus, we have
+$\begin{aligned} \frac{dG(\eta)}{d\eta}=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2},\end{aligned}$
 and
-$\frac{d^{2}G(\eta)}{d\eta^{2}}=L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\sigma^{2}.$
+$\begin{aligned} \frac{d^{2}G(\eta)}{d\eta^{2}}=L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\sigma^{2}.\end{aligned}$
 Since $L_{1}$, $E$, and $\sigma^{2}$ are all greater than 0, we have $\frac{d^{2}G(\eta)}{d\eta^{2}}\textgreater 0.$ Thus, $G(\eta)$ is proved to be a convex function and there exists a minimum value of $G(\eta)$.
@@
 Proof.
-When $G^{\prime}(\eta)$ equals 0, then $G(\eta)$ obtains a extremum value. Let $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=0,$ so we get
+When $G^{\prime}(\eta)$ equals 0, then $G(\eta)$ obtains a extremum value. Let $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=0$ , we can get:
 $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2}=0.$
-Thus, $G(\eta)$ exists a extremum value when $\eta=\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}.$ Because corollary 1 proves the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative, the extremum value is the minimal value.
-Therefore, we can understand that the variation of loss function exists a minimum bound when $\eta=\eta^{\star}$.
+Thus, $G(\eta)$ exists a extremum value when $\eta=\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}$ .
+Because corollary 1 proves the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative, the extremum value is the minimal value.
+Therefore, we can understand that it exists a minimum bound when $\eta=\eta^{\star}$.
 ∎
@@
 Proof.
-To guarantee that the local loss function decreases after each communication, we need to make sure that $G(\lambda)\textless 0$, so we have
-$-\left(\eta-\frac{L_{1}\eta^{2}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda\textless 0.$
-After simplification, we get
-$\lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.$
+To guarantee that the local loss function decreases after each communication, it is necessary to make sure that $G(\lambda)\textless 0$. We have:
+$\begin{aligned} -\left(\eta-\frac{L_{1}\eta^{2}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda\textless 0.\end{aligned}$
+After simplification, we can get:
+$\begin{aligned} \lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.\end{aligned}$
+Theorem 2.
+Let the proportions of benign and malicious clients be $(1-\kappa)$ and $\kappa$, respectively, with $\kappa\in[0,1]$.
+In the $t$-th round, let $\textit{{u}}_{ben}^{k}$ denote the average prototype of benign clients for class $k$, and $\textit{{u}}_{mal}^{k}$ denote the average prototype of malicious clients for class $k$.
+Then, the aggregated global prototype for class $k$ is given by:
+$\textit{{C}}^{k}_{t+1}=(1-\kappa)\textit{{u}}_{ben}^{k}+\kappa\textit{{u}}_{mal}^{k}$
+and its deviation in magnitude from the benign prototypes satisfies:
+$\|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}=\kappa\|\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}\|_{2}\leq 2\kappa.$
+From the above formula, we can observe that the deviation of the global prototype from the benign prototype increases linearly with the malicious proportion $\kappa$, thus ensuring that the influence of malicious clients always remains within a controllable range.
+Proof.
+By definition of the malicious client proportion $\kappa$, the global prototype can be expressed as:
+[…]
+Subtracting $\textit{{u}}_{ben}^{k}$ from both sides yields:
+$\begin{aligned} \textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}=\kappa(\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}).\end{aligned}$
+Taking the $\ell_{2}$-norm on both sides, we have:
+$\begin{aligned} \|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}=\kappa\|(\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k})\|_{2}.\end{aligned}$
+Since $\textit{{u}}_{mal}^{k}$ and $\textit{{u}}_{ben}^{k}$ are unit vectors, the triangle inequality gives:
+$\begin{aligned} \|\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}\|_{2}\leq\|\textit{{u}}_{mal}^{k}\|+\|\textit{{u}}_{ben}^{k}\|=2.\end{aligned}$
+Consequently, we obtain the relation $\|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}\leq 2\kappa.\$This completes the proof of Theorem 2. ∎
 Theorem 3.
@@
 Proof.
-During the secure aggregation, the two servers can obtain plaintext information $\|\textit{{C}}^{\prime k}_{t+1}\|$ and $\textit{Sum}^{k}_{t}$, where $\|\textit{{C}}^{\prime k}_{t+1}\|$ denotes the norm of trusted update, $\textit{Sum}^{k}_{t}$ denotes the sum of aggregation weight of class $k$.
+During the secure aggregation, the two servers can obtain plaintext values $\|\textit{{C}}^{\prime k}_{t+1}\|$ and $\textit{Sum}^{k}_{t}$, where $\|\textit{{C}}^{\prime k}_{t+1}\|$ denotes the norm of trusted prototype, $\textit{Sum}^{k}_{t}$ denotes the sum of aggregation weight of class $k$.
 For non-colluding Aggregator and Verifier, they cannot get any sensitive information from these plaintext information.
-In addition, for colluding malicious clients, when there are $(|\mathcal{S}|-1)$ malicious clients in PPFPL, they can infer the local prototype about benign client from the encrypted global prototype.
+In addition, for colluding malicious clients, if $(|\mathcal{S}|-1)$ clients are compromised, they can theoretically infer the local prototype of the remaining benign from the encrypted global prototype.
 However, the real scenario does not exist when there are $(|\mathcal{S}|-1)$ malicious clients.
@@
 ∎
-VII Experiments
+7 Experiments
 In this section, we evaluate the performance of PPFPL in the presence of data poisoning attacks on Non-IID data.
-VII-A Experimental Settings
-VII-A1 Datasets and Models
-Similar to previous works [16][36][52], we utilize three public available datasets, namely MNIST, FMNIST, and CIFAR10, to evaluate performance of our PPFPL.
-  1. 1.
-    MNIST: This dataset comprises handwritten digits, with a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each represented as a 28 $\times$ 28 gray-scale image.
-  2. 2.
-    FMNIST: The FMNIST is a clothing image dataset containing a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each of which is a 28$\times$28 grey-scale image.
-  3. 3.
-    CIFAR10: As a color image dataset, CIFAR10 contains a training set of 50,000 samples in 10 classes and a test set of 10,000 samples, with each image measuring 32$\times$32 pixels.
+7.1 Experimental Settings
+7.1.1 Datasets and Models
+Similar to previous works [16][38][55], we utilize three public available datasets, namely MNIST [56], FMNIST [57], and CIFAR10 [58], to evaluate performance of our PPFPL.
 Furthermore, we apply Convolution Neural Network (CNN) as local model to both MNIST and FMNIST.
-For CIFAR-10, we employ ResNet18 as the local model, initialized with pre-trained parameters.
+For CIFAR10, we employ ResNet18 as the local model, initialized with pre-trained parameters.
 These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
-VII-A2 Hyper-parameters Settings of FL
-We employ a cross-silo configuration in FL setup.
+7.1.2 Hyper-parameters Settings of FL
+We employ a cross-silo configuration in our FL experiments.
 Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round.
-The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively.
-We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to $E=5$.
+The number of rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively.
+The local learning rate is set to $\eta=0.01$, the importance weight to $\lambda=1$, the batch size to 64, and the number of local iterations to $E=5$.
+The detection threshold $\chi$ is configured as 0.
 Except for special declared, the above default settings are applied.
 These hyper-parameters are consistent across all clients.
-VII-A3 Non-IID Settings
-To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios.
-Specifically, large organizations (e.g., hospitals, companies) possess different data classes, and their class distributions may differ significantly, or even be missing some classes altogether.
+For encryption, we instantiate the CKKS in our experiments using TenSEAL [59] with the polynomial modulus degree of 8192 and a coefficient modulus of 200-bits, and a scale of 40-bits.
+This configuration offers 128-bits security level [60][61], and all encryption parameters follow the recommended settings provided by TenSEAL.
+7.1.3 Non-IID Settings
+To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in cross-silo scenarios.
+Specifically, large organizations (e.g., hospitals, companies) hold datasets with different classes, and their class distributions may differ significantly, or even be missing some classes altogether.
 When these organizations participate in federated training, the union of their data classes defines the entire FL classification task.
 This phenomenon leads to Non-IID data across organizations.
-The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a positive percentage of every class.
+However, the Dirichlet distribution [62] assumes that each client’s data is sampled from all classes, meaning each client usually has samples of every class.
 However, in cross-silo scenarios, some clients may have no samples from certain classes, which contradicts the assumptions of Dirichlet distribution.
 To model the data distribution in cross-silo scenarios, similar to previous works [23][63], we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts.
-In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity.
-Clients are randomly assigned classes, with partial class overlap among them.
+In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, to create the class-space heterogeneity.
+Clients are randomly assigned to different classes, with partial class overlap among them.
 To visualize the different data distributions, we plot heat maps as shown in Fig.
@@
 ((c)) Avg=5, Std=2
-Figure 2: Heat maps for different data distributions under heterogeneity distribution.
-VII-A4 Setting of Data Poisoning Attacks
-In our experiments, we consider two types of data poisoning attacks: feature attacks and label attacks.
+Figure 2: Heat maps under different heterogeneity distributions.
+7.1.4 Setting of Data Poisoning Attacks
+In our experiments, we evaluate two types of data poisoning attacks: feature attacks and label attacks.
 For feature attacks, malicious clients randomly alter their own training data features in a completely randomized manner without following any specific rule.
-For label attacks, malicious clients alter the labels of their own training data to incorrect labels.
+For label attacks, malicious clients modify the labels of their training data to incorrect labels.
 All training data of malicious clients are modified.
-Malicious clients do not dynamically change their behaviors during federated training, and the attack remains consistent throughout the training period.
-Since the number of malicious clients affects FL performance differently, we set different proportions of malicious clients.
-Formally, Att is set as the proportion of malicious clients.
-VII-A5 Evaluation Measure
-Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks.
-Therefore, we evaluate the FL performance by testing the average accuracy of benign clients.
-For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks.
-In addition, we compare our PPFPL with robust schemes (i.e., Krum [35], Foolsgold [54], and FedDMC [36]) and privacy-preserving robust schemes (i.e., ShieldFL [18], PBFL [16], and AntidoteFL [20]) under the same experiment conditions.
-VII-B Experimental Results
-VII-B1 Visualization of Prototypes
-We visualize the prototypes submitted by clients in PPFPL using t-SNE, as shown in Fig.
+Unless otherwise specified, malicious clients keep their behaviors fixed during federated training, and the attack remains consistent throughout the training period.
+Since the number of malicious clients affects FL performance differently, we set different proportions of malicious clients, denotes as Att.
+7.1.5 Evaluation Measure
+Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while maintaining robustness against data poisoning attacks.
+Therefore, we evaluate the FL performance by measuring the average accuracy of benign clients.
+For experiment comparison, we use FedAvg as the baseline, assuming it is not subject to data poisoning attacks.
+In addition, we compare PPFPL with robust schemes (i.e., Krum [37], Foolsgold [64], and FedDMC [38]) and privacy-preserving robust schemes (i.e., ShieldFL [19], PBFL [16], AntidoteFL [21], and FLDP-DM [65]) under the same experiment conditions.
+7.2 Experimental Results
+7.2.1 Visualization of Prototypes
+We visualize client-submitted prototypes in PPFPL using t-SNE, as shown in Fig.
 3.
-Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on the CIFAR10 dataset within a single communication round.
-We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks.
+Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on CIFAR10 within a single communication round.
+We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of features attacks or labels attacks.
 In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area.
-This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas.
+This is because when malicious clients perform local model training, the features or labels are disrupted, causing the generated prototypes to spread from the center point to the surrounding areas.
 Most importantly, we observe a clear difference between the prototypes submitted by benign and malicious clients.
 This difference is prominent because it is independent of the data distribution.
-The finding validates the effectiveness of using prototypes to defend against data poisoning attacks.
-To filter out these malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results.
+The finding validates the effectiveness of leveraging prototypes to defend against data poisoning attacks.
+To effectively filter out malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results.
 The following is the security evaluation of PPFPL.
-((a)) Feature attack,$\textit{Avg}=3,\textit{Std}=2$.
-((b)) Feature attack,$\textit{Avg}=5,\textit{Std}=1$.
-((c)) Label attacks, $\textit{Avg}=3$, $\textit{Std}=2$.
-((d)) Label attacks, $\textit{Avg}=5$, $\textit{Std}=1$.
-Figure 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
-VII-B2 Security Evaluation
+((a)) Feature attacks, Avg=3,Std=2.
+((b)) Feature attacks, Avg=5,Std=1.
+((c)) Label attacks, Avg=3,Std=2.
+((d)) Label attacks, Avg=5,Std=1.
+Figure 3: The t-SNE visualization of client-submitted prototypes.
+Different colors indicate distinct classes.
+The symbol $\cdot$” denotes prototypes from benign clients, and $\times$” represents those from malicious clients.
+7.2.2 Security Evaluation
 ((a)) Feature attacks, FMNIST.
@@
 Figure 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
-((a)) Feature attacks, CIFAR10.
-((b)) Label attacks, CIFAR10.
-Figure 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
-To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$.
-Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$.
-Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, and AntidoteFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
+To evaluate the security of our framework, we evaluate PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with Avg=3 and Std=2.
+Specifically, the proportion of malicious clients is set to Att=20%.
+Furthermore, we compare PPFPL with existing defense schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, AntidoteFL, and FLDP-DM), where the global model learned by FedAvg is not subject to data poisoning attacks.
 The experiment results are shown in Fig.
 4, where the black dashed line represents the highest accuracy of PPFPL without attacks.
-We observe that the accuracy of ShieldFL and PBFL (under attacks) is severely lower than FedAvg (no attack) in CIFAR10.
-This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious.
-This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance.
-Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of the model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning.
-Moreover, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence.
+We observe that ShieldFL and PBFL suffer significant accuracy degradation under attacks compared to FedAvg without attacks on the CIFAR10 dataset.
+This indicates that as datasets become more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious.
+This is mainly because complex datasets increase the difficulty of defense, and poisoned Non-IID data further degrades the defense performance.
+Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning.
+Moreover, PPFPL (under attacks) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence.
 This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients.
@@
 5.
-We notice that the loss of PBFL fluctuates significantly during training process, while PPFPL is relatively smooth.
-This indicates that data poisoning attacks disrupt the convergence of PBFL, while the convergence of PPFPL is not affected by data poisoning attacks, thus ensuring the accuracy of the model.
+We observe that PBFL’s loss fluctuates significantly throughout the training process, while PPFPL demonstrates relatively smooth convergence.
+This indicates that data poisoning attacks severely disrupt PBFL’s convergence, while the convergence of PPFPL is not affected by data poisoning attacks, thereby preserving model accuracy.
 This is consistent with our Corollary 3.
-Specifically, PPFPL still satisfies convergence in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them.
+Specifically, PPFPL still converges in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them.
 Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security.
-Notably, our experiments show that the performance impact of feature attacks and label attacks is similar for PPFPL and other schemes.
+Notably, our experiments show that feature attacks and label attacks have similar impacts on the performance of PPFPL and other schemes.
 This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
-((a)) Feature attacks
-((b)) Label attacks
-Figure 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
-VII-B3 Different Data Distributions for PPFPL
-To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10.
-In addition, we compare the performance of Krum, Foolsgold, ShieldFL, PBFL, FedDMC, and AntidoteFL.
-The specific experimental results are reported in Table I.
-Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across the communication rounds.
-From Table I, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution.
+((a)) Feature attacks, CIFAR10.
+((b)) Label attacks, CIFAR10.
+Figure 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
+7.2.3 Different Data Distributions for PPFPL
+To further test the security of PPFPL under different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10.
+In addition, we compare its performance with Krum, Foolsgold, ShieldFL, PBFL, FedDMC, AntidoteFL, and FLDP-DM.
+The specific experimental results are reported in TABLE 2.
+Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across all communication rounds.
+From TABLE 2, we can observe that the performance of PPFPL is not easily affected by changes in data distribution, while other schemes are very susceptible to data distribution.
 This is attributed to that the client-submitted prototype does not change due to the change in data distribution.
@@
 This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon.
-Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
-TABLE I: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
+Overall, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, making it suitable for real-world deployments.
+Table 2: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
 Dataset Method Att % Std Test Average Accuracy (%)
 Avg = 3 Avg = 4 Avg = 5 Avg = 6 Avg = 7
-MNIST Krum 20 1 95.96 96.95 97.01 97.04 97.12
+MNIST Krum [37] 20 1 95.96 96.95 97.01 97.04 97.12
 2 95.67 94.28 94.88 95.51 96.20
@@
 2 95.67 94.28 94.88 94.77 95.80
-Foolsgold 20 1 96.69 94.03 95.13 95.76 95.39
+Foolsgold [64] 20 1 96.69 94.03 95.13 95.76 95.39
 2 96.19 94.82 96.73 95.27 95.75
@@
 2 96.02 94.74 96.58 95.01 95.49
-PBFL 20 1 97.78 97.57 96.34 97.10 96.98
+PBFL [16] 20 1 97.78 97.57 96.34 97.10 96.98
 2 97.36 96.17 96.26 94.64 95.41
@@
 2 95.40 93.62 93.48 94.70 94.62
-ShieldFL 20 1 96.91 94.05 95.10 95.64 95.25
+ShieldFL [19] 20 1 96.91 94.05 95.10 95.64 95.25
 2 97.25 94.88 95.69 95.44 95.76
@@
 2 97.08 94.69 95.46 95.73 94.46
-FedDMC 20 1 97.62 97.46 97.16 96.86 96.53
+FedDMC [38] 20 1 97.62 97.46 97.16 96.86 96.53
 2 97.53 97.32 97.04 96.73 96.39
@@
 2 97.43 96.58 96.34 96.26 94.33
-AntidoteFL 20 1 97.58 97.40 97.09 96.73 96.34
+AntidoteFL [21] 20 1 97.58 97.40 97.09 96.73 96.34
 2 97.31 97.18 96.85 95.45 96.30
@@
 2 97.08 94.69 95.46 95.73 94.46
+FLDP-DM [65] 20 1 97.60 97.22 97.15 96.78 96.37
+2 97.48 97.16 96.93 96.40 96.08
+30 1 97.14 96.85 96.77 96.76 96.42
+2 96.95 96.36 96.12 95.91 94.87
 PPFPL 20 1 97.87 97.53 97.35 97.08 96.89
@@
 2 97.52 96.92 96.70 96.50 96.48
-FMNIST Krum 20 1 84.24 84.80 86.03 86.52 87.35
+FMNIST Krum [37] 20 1 84.24 84.80 86.03 86.52 87.35
 2 85.02 83.64 86.82 84.75 86.08
@@
 2 84.62 82.73 85.73 82.86 84.09
-Foolsgold 20 1 84.88 84.73 84.04 83.55 82.62
+Foolsgold [64] 20 1 84.88 84.73 84.04 83.55 82.62
 2 82.49 83.88 85.31 81.75 82.24
@@
 2 81.86 83.43 84.98 81.19 82.68
-PBFL 20 1 87.01 80.37 76.93 74.26 72.10
+PBFL [16] 20 1 87.01 80.37 76.93 74.26 72.10
 2 87.42 76.31 74.47 72.99 72.43
@@
 2 79.34 72.68 72.81 70.57 70.49
-ShieldFL 20 1 87.01 78.90 78.06 70.90 70.26
+ShieldFL [19] 20 1 87.01 78.90 78.06 70.90 70.26
 2 76.47 77.89 70.96 67.47 66.80
@@
 2 75.81 77.35 70.57 66.42 65.46
-FedDMC 20 1 90.60 88.21 87.03 86.40 86.11
+FedDMC [38] 20 1 90.60 88.21 87.03 86.40 86.11
 2 89.53 88.70 87.27 87.20 87.13
@@
 2 89.40 87.52 86.23 86.40 86.12
-AntidoteFL 20 1 89.13 87.20 87.03 86.40 86.11
+AntidoteFL [21] 20 1 89.13 87.20 86.58 85.36 85.25
 2 88.72 86.13 85.34 85.16 86.48
@@
 2 87.10 86.61 85.10 84.39 84.90
+FLDP-DM [65] 20 1 90.48 88.06 86.53 86.17 85.86
+2 89.32 87.94 86.95 85.09 84.77
+30 1 88.64 87.35 85.90 85.47 85.30
+2 88.79 86.92 85.71 85.06 84.55
 PPFPL 20 1 91.38 90.18 89.27 88.56 88.13
@@
 2 90.40 88.62 87.74 88.41 87.97
-CIFAR10 Krum 20 1 62.91 61.24 68.19 68.87 69.22
+CIFAR10 Krum [37] 20 1 62.91 61.24 68.19 68.87 69.22
 2 57.90 57.94 68.33 68.65 68.79
@@
 2 56.67 57.29 67.03 68.49 68.51
-Foolsgold 20 1 60.91 59.24 66.19 66.45 66.52
+Foolsgold [64] 20 1 60.91 59.24 66.19 66.45 66.52
 2 56.90 55.94 66.33 66.62 66.70
@@
 2 56.17 55.09 65.85 66.40 66.41
-PBFL 20 1 38.06 39.18 41.81 42.42 42.75
+PBFL [16] 20 1 38.06 39.18 41.81 42.42 42.75
 2 39.46 40.54 41.30 42.12 42.76
@@
 2 37.80 37.52 41.08 41.96 42.67
-ShieldFL 20 1 61.62 61.94 68.31 69.53 68.66
+ShieldFL [19] 20 1 61.62 61.94 68.31 69.53 68.66
 2 57.84 57.26 66.14 67.43 68.62
@@
 2 57.10 56.76 65.47 66.42 67.69
-FedDMC 20 1 75.28 74.16 74.37 73.86 73.50
+FedDMC [38] 20 1 75.28 74.16 74.37 73.86 73.50
 2 73.52 72.40 72.37 72.06 71.58
@@
 2 72.67 72.48 72.11 71.85 70.78
-AntidoteFL 20 1 72.66 72.27 71.19 72.19 70.65
+AntidoteFL [21] 20 1 72.66 72.27 71.19 72.19 70.65
 2 70.86 70.40 70.19 69.84 69.75
@@
 2 69.82 68.70 68.42 68.83 68.17
+FLDP-DM [65] 20 1 74.89 72.48 72.07 70.63 68.30
+2 72.30 71.55 70.50 69.38 69.60
+30 1 73.86 71.79 70.49 70.43 68.02
+2 72.05 71.33 69.54 69.26 67.40
 PPFPL 20 1 83.41 80.95 77.82 74.76 71.35
@@
 2 83.16 82.37 77.47 74.62 71.19
-VII-B4 Secure Aggregation Protocol for PPFPL
-To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
-In addition, the $\boldsymbol{X}$ in Table II denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging.
-As shown in Table II, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings.
-The performance degradation stems from the influence of malicious clients’ prototypes on the aggregation process, leading to a poisoned global prototype.
+7.2.4 Secure Aggregation Protocol for PPFPL
+To test the effectiveness of the aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively.
+Here, $\chi=-1$ means that the protocol performs normalization verification and average aggregation.
+The data distribution is set to Non-IID (i.e., Avg=3, Std=2).
+In addition, $\boldsymbol{X}$ in Table 3 denotes that client-submitted prototypes are unnormalized and aggregated using simple averaging.
+As shown in TABLE 3, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings.
+The performance degradation is because malicious prototype poisons the global prototype.
 However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss.
-Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients.
-Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol.
+Thus, the model accuracy among benign clients in PPFPL remains relatively robust against the influence of malicious clients.
+Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results.
 The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks?
-We explore this question in the next experiment.
-TABLE II: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
+We address this question in the subsequent experiments.
+Table 3: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
 Datasets Att $\chi$ = -1 $\chi$ = 0 $\chi$ = 0.2 $\chi$ = 0.5 $\boldsymbol{X}$
@@
 Figure 6: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
-VII-B5 High Proportion Attacks for PPFPL
-To test the performance of PPFPL under high proportion attacks, we evaluate it with Att = 20%, 40%, 60% and 80%, respectively.
-In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
-The $\lambda$ denotes the importance weight of the auxiliary term in loss function, which can be considered as the degree of influence among clients.
+7.2.5 High Proportion Attacks for PPFPL
+To evaluate the performance of PPFPL under high proportion attacks, we test it with Att = 20%, 40%, 60% and 80%.
+In addition, we set $\lambda$ as 0.01, 0.1 and 1 under Non-IID data distribution (Avg=3, Std=2).
+The $\lambda$ denotes the importance weight of the auxiliary term in the loss function, which can be considered as the degree of influence among clients.
 The experimental results are shown in Fig.
-7.
-We surprisingly observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients.
-This is because the benign client does not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data.
+6.
+Surprisingly, We observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients.
+This is because the benign clients do not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data.
 Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets.
@@
 This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients.
-Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but increases resistance to data poisoning attacks.
+Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but enhances resistance to data poisoning attacks.
 Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
-VII-B6 Stability of PPFPL
+7.2.6 Dynamic Data Poisoning Attacks for PPFPL
+In practical cross-silo scenarios, the behavior of malicious clients may vary across different rounds.
+To emulate this scenario, we configure each malicious client to alternate between feature attacks and label attacks in consecutive rounds, termed dynamic poisoning attacks.
+We evaluate the performance of PPFPL under the dynamic attacks.
+The experimental results (shown in TABLE 4) demonstrate that the performance of PPFPL only slightly decreases with the increase in the number of malicious clients, mainly due to the reduction in benign samples.
+Thus, PPFPL maintains superior performance under dynamic poisoning attacks.
+This is because, regardless of how malicious clients manipulate their submitted prototypes, our secure aggregation protocol can limit their impact to the directional perturbation.
+The protocol filters out prototypes with excessive directional deviation, ensuring that PPFPL is resilient to dynamic poisoning attacks.
+Table 4: Test Average Accuracy (%) of PPFPL against dynamic poisoning attacks.
+Datasets
+Data
+Distributions
+Proportion of Malicious Clients (%)
+0% 10% 20% 30% 40%
+CIFAR10 Avg=3, Avg=1 83.05 83.20 82.79 82.27 81.55
+Avg=3, Avg=2 83.29 83.13 82.83 82.04 81.26
+Avg=4, Avg=1 82.04 81.87 80.71 80.93 80.60
+Avg=4, Avg=2 82.11 82.06 82.45 81.75 81.21
+Avg=5, Avg=1 77.82 77.31 77.24 76.92 76.49
+Avg=5, Avg=2 77.95 78.12 78.03 77.80 76.22
+7.2.7 Stability of PPFPL
 ((a)) Feature attacks
@@
 Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments.
-To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious.
-The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$).
+To verify the stability of PPFPL, we conduct experiments on MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious.
+The data distribution follows the Non-IID setting (Avg=3, Std=2).
 Each experiment is repeated 10 times.
-The utilization of boxplots in Fig.
+The boxplots in Fig.
 7 provides a clear visualization of the stability of PPFPL.
@@
 The stability of PPFPL proves its reliability in real-world scenarios.
-VII-B7 Scalability of PPFPL
-Federated learning is often deployed with an uncertain number of clients, and scalability directly determines whether PPFPL can maintain stable performance in environments with different numbers of clients.
-To evaluate the scalability of PPFPL, we conducted experiments on the MNIST, FMNIST, and CIFAR10 datasets with 20% malicious clients, and the number of clients set to 20, 25, 30, 35, and 40.
-The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=1$).
+7.2.8 Scalability of PPFPL
+Federated learning is often deployed on an unknown number of clients.
+Scalability directly determines whether PPFPL can maintain stable performance under different client scales.
+To evaluate the scalability of PPFPL, we conduct experiments on MNIST, FMNIST, and CIFAR10 datasets with 20% malicious clients, varying the number of clients as 20, 30, 40, 80, and 100.
+The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=1$) setting.
 Fig.
 8 shows the scalability of PPFPL against data poisoning attacks.
-It is observed that the performance of PPFPL drops by 0.24% on MNIST and 0.42% on CIFAR10 when the number of participants is increased by 100%, which is normal in federated learning.
-Thus, PPFPL shows excellent scalability against data poisoning attacks with data heterogeneity.
+It is observed that PPFPL maintains stable performance when the number of clients is 40.
+However, the performance of PPFPL degrades significantly when the number increases to 100.
+For example, the accuracy drops by 1.72% on MNIST and 14.11% on CIFAR10.
+The main reason is that PPFPL follows a personalized training paradigm in which each client relies on its own data for training.
+As the number of clients increases, the amount of data allocated to each client decreases.
+The small amount of data in each client makes it difficult to obtain a high-quality model through local training.
+This shows that the performance of PPFPL is closely related to the number of clients and ultimately depends on on the data available to each client.
+Therefore, in large-scale networks, each client needs to have sufficient data to fully exploit the advantages of PPFPL.
+Under our experimental conditions, the data volumes of MNIST, FMNIST, and CIFAR10 datasets can reasonably support up to about 40 clients, within which PPFPL demonstrates good scalability.
 ((a)) MNIST
@@
 Figure 8: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
-VII-B8 Impact of Local Iterations
-We investigate the effect of the number of local iterations on the test average accuracy for PPFPL.
-We conduct experiments on MNIST, FMNIST, and CIFAR10 datasets, where 20% of malicious clients launched feature attacks.
-The number of local iterations are 1, 5, 10, 15, 20 and 25.
-The results are shown in Fig.
-10.
-We observe that when the number of local iterations is 1, the changes in the local model are very small, resulting in slow training and relatively low accuracy for the same number of communication rounds.
-When the number of local iterations is 5, PPFPL has the highest test average accuracy on the MNIST dataset.
-However, on CIFAR10, the optimal number of local iterations should be set to 15.
-In addition, when the number of local iterations is too high, the test average accuracy decreases, which is due to the overfitting of the local model.
-((a)) MNIST.
-((b)) FMNIST.
-((c)) CIFAR10.
-Figure 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
-VII-B9 Efficiency Evaluation
+7.2.9 Ablation Study
+To comprehensively evaluate the effectiveness of the secure aggregation protocol, we conduct ablation experiments.
+Within the PPFPL framework, we remove the normalization verification module (denoted as “w/o NV”, where “w/o” stands for “without”) and the secure aggregation protocol (denoted as “w/o SAP”).
+We conduct experiments on CIFAR10 dataset, where 30% of malicious clients launched attacks.
+The experimental results are shown in TABLE 5.
+Under both attack settings, PPFPL’s performance degrades regardless of which component is removed, demonstrating the effectiveness of normalization verification and the secure aggregation protocol.
+Notably, in the absence of malicious clients, removing the secure aggregation protocol actually led to optimal model performance.
+This phenomenon is attributed to the increase in the number of benign training samples, which enhances the overall FL performance.
+Table 5: Performance impact of different components in PPFPL.
+  Schemes
+  Data
+  Distribution
+  Test Average Accuracy (%)
+  Feature   Label   No
+  PPFPL   Avg=3, Avg=1   82.31   82.50   83.07
+  Avg=3, Avg=2   83.16   83.42   83.29
+  w/o NV   Avg=3, Avg=1   80.26   80.29   83.20
+  Avg=3, Avg=2   80.74   80.61   83.34
+  w/o SAP   Avg=3, Avg=1   77.69   77.24   83.53
+  Avg=3, Avg=2   78.03   78.53   83.64
+Note: “Feature” indicates feature attacks, “Label” indicates label attacks, and “No” indicates the absence of malicious clients.
+In addition, “NV” indicates the normalization verification module, “SAP” indicates the secure aggregation protocol.
+Best rusults are in blod.
+7.2.10 Efficiency Evaluation
 ((a))
@@
 (b) The number of parameters submitted by client using CNN model in one round under different scheme.
-To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as shown in Fig.
-11.
-We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same amount of parameters.
-Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture.
-This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, rather than the model architecture.
-Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
-TABLE III: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
+To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as reported in Fig.
+9.
+We observe that both PBFL and ShieldFL require clients to submit a fixed number of parameters determined by the model architecture, resulting in identical parameter volumes across all clients.
+In contrast, the number of parameters submitted by clients in PPFPL is significantly lower under the same model architecture.
+This is because the number of prototype parameters in PPFPL is determined by the output of the feature extractor rather than the model architecture itself.
+Consequently, PPFPL effectively reduces the parameter volume transmitted to the servers, thereby decreasing both computational and communication overhead.
+Table 6: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
    Operation    PBFL    ShieldFL    PPFPL
@@
    Decrypt    1.35s    1.19s    0.40s
-In addition, ciphertext operations are the primary factor affecting the efficiency of our framework.
-We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations.
-Note that PBFL is a federated learning scheme based on CKKS, while ShieldFL is built upon the two-door HE.
-For each client, we use the time spent on encryption and decryption for each iteration as a metric.
-As shown in Table III, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL.
-This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
-VII-C Complexity Analysis
-TABLE IV: Computation Overload and Communication Overload of client.
+Furthermore, ciphertext operations are the primary factor affecting the efficiency of our framework.
+We evaluate the time cost of ciphertext operations for PPFPL, PBFL, and ShieldFL.
+Notably, PBFL employs the CKKS, while ShieldFL is built upon two-door HE.
+We measure the time spent on encryption and decryption per round for each client as our evaluation metric.
+As shown in Table 6, PPFPL’s performance in both encryption and decryption is lower than that of PBFL and ShieldFL.
+This advantage stems from the fact that PPFPL requires fewer prototype parameters submitted by each client, significantly reducing the computational cost of the ciphertext operations.
+7.3 Complexity Analysis
+Table 7: Computation overload and communication overload of client.
 Scheme Computation Overload Communication Overload
@@
 PPFPL $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ $\mathcal{O}(nP_{h})$
-We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15].
-The results are shown in Table IV.
-The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption.
-Formally, the computational overhead is expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ denotes the prototype parameters, $T_{ch}$ denotes the time overload of encryption, $T_{tr}$ denotes the time overhead of local model training, and $T_{pro}$ denotes the time overhead of prototype generation.
+We analyze the computational and communication overload for each client in PPFPL and compared it with similar schemes [16][19][15], as shown in TABLE 7.
+The computational overload for clients in PPFPL consists of three components: local model training, prototype generation, and encryption.
+Formally, the computational overhead can be expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ is the number of prototype parameters, $T_{ch}$ denotes the encryption time overhead, $T_{tr}$ denotes the time overhead for local model training, and $T_{pro}$ denotes the time overhead of prototype generation.
 Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$.
-Other similar schemes [16][18][15] adopt homomorphic encryption on gradients, their encryption overhead is $\mathcal{O}(gT_{ch})$, where $g$ denotes the gradient parameters.
-Since $g>p$, the overload satisfies $\mathcal{O}(gT_{ch})>\mathcal{O}(pT_{ch})$.
-Hence, the relationship $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ holds.
-Consequently, PPFPL imposes a reasonable computational overhead on clients.
-In addition, the client communication overhead is $\mathcal{O}(nP_{h})$ in PPFPL, where $P_{h}$ denotes the communication complexity of a number.
-Compared with [16][18][15], the relationship $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds.
-VIII Concludes
-In this paper, we propose PPFPL, a privacy-preserving federated prototype learning framework, which improves performance while resisting data poisoning attacks in poisoned Non-IID data.
-Our work is the first to introduce prototype learning into PPFL to address privacy threats and data poisoning attacks in across-silo FL.
-As for how to implement lightweight privacy-preserving FL to resist data poisoning attacks on Non-IID data, we leave them to future exploration.
+Other comparable schemes [16][19][15] adopt homomorphic encryption on gradients, resulting in an encryption overhead $\mathcal{O}(gT_{ch})$, where $g$ denotes the number of gradient parameters.
+Since the relation $g\textgreater p$ holds, the overload relation satisfies $\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(pT_{ch})$.
+Hence, the relation of ciphertext operations is $\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(pT_{ch})$.
+In addition, the client communication overhead in PPFPL is $\mathcal{O}(nP_{h})$, where $P_{h}$ denotes the communication complexity per parameter.
+Compared with the schemes in [16][19][15], the inequality $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds, demonstrating the efficiency advantage of PPFPL.
+8 Conclusion
+This paper proposes PPFPL, a privacy-preserving federated prototype learning framework that effectively enhances federated learning performance in poisoned Non-IID data environments.
+Through comprehensively theoretical analysis and experimental validation, we demonstrate the significant advantages of PPFPL compared existing FL defense schemes.
+The framework provides robust security guarantees for distributed computing scenarios, with important practical applications in privacy-sensitive domains such as finance and healthcare.
+However, our current framework has certain limitations.
+First, frequent interactions in the two-server architecture increase system failures caused by communication latency.
+Second, existing experimental datasets do not adequately capture the complexity of real-world cross-silo scenarios, making it difficult to effectively evaluate prototypes generated by benign clients under varying data quality conditions.
+This limitation may result in unfair evaluation outcomes in practical settings.
+Therefore, our future work will focus on two directions: (i) designing more efficient two-server interaction protocols while ensuring privacy and security.
+(ii) constructing benchmark datasets that capture the characteristics of cross-silo scenarios to evaluate FL system performance.